### v2

# The Best Instruction-Tuning Data are Those That Fit

Abstract
High-quality supervised finetuning (SFT) data are crucial for eliciting strong capabilities from pretrained large language models (LLMs). Typically, instructions are paired with multiple responses sampled from other LLMs, which are often out of the distribution of the target model to be finetuned. This, at scale, can lead to diminishing returns and even hurt the models’ performance and robustness. We propose GRAPE, a novel SFT framework that accounts for the unique characteristics of the target model. For each instruction, it gathers responses from various LLMs, and selects the one with the highest probability measured by the target model, indicating that it aligns most closely to the target model’s pretrained distribution; it then proceeds with standard SFT training. We first evaluate GRAPE with a controlled experiment, where we sample various solutions for each question in UltraInteract from multiple models and finetune commonly used LMs like Llama3.1-8B, Mistral-7B and Qwen2.5-7B on GRAPE-selected data. GRAPE significantly outperforms strong baselines, including distilling from the strongest model with absolute gain up to 13.8% averaging across benchmarks, and training on 3$\times$ more data with maximum 17.3% performance improvements. GRAPE’s strong performance generalizes to realistic settings. We experiment with the post-training data used for Tulu3 and Olmo-2. GRAPE can outperform strong baselines with 4.5 times the data by 6.1% and state-of-the-art data selection approach by 3.9% on average performance. Remarkably, using 1/3 data and half number of epochs, GRAPE allows Llama3.1-8B to surpass the performance of Tulu3-SFT by 3.5%.
Machine Learning, ICML
1 Introduction
High-quality, large-scale supervised data is crucial for supervised instruction finetuning (SFT; Databricks, 2023; Köpf et al., 2023; Zhao et al., 2024a; Zheng et al., 2024). A common practice of data collection involves sampling responses from strong language models, predominantly focusing on expanding the size of the dataset and improving the overall quality of the responses (Sun et al., 2023; Taori et al., 2023; Wang et al., 2023; Xu et al., 2024c; Chen et al., 2024). However, recent research suggests that there is more complex dynamics involved (Xu et al., 2024d). A plateau effect in synthetic data scaling, where performance either stagnates or even declines as the size of the synthetic data increases beyond a certain point, has been widely observed. This phenomenon arises due to issues such as diminishing diversity (Padmakumar & He, 2024; Guo et al., 2023) and distortion in the data distribution (LeBrun et al., 2021), which ultimately undermine the base model’s performance and robustness (Alemohammad et al., 2024; Gerstgrasser et al., 2024; Shumailov et al., 2023; Dohmatob et al., 2024; Hataya et al., 2023; Martínez et al., 2023a, b; Bohacek & Farid, 2023; Briesch et al., 2023).
Thus, effective instruction tuning requires more than scaling up the data; it often needs “tailoring” the data to the unique characteristics of the target model. Existing works focus on enhancing the model’s existing knowledge and capabilities (Du et al., 2023) and optimizing the curriculum progression for instruction tuning (Zhao et al., 2024b; Lee et al., 2024a; Feng et al., 2023; Setlur et al., 2024). They typically prescribe questions for each model to learn solutions. However, it remains elusive which responses best suit a given base model for fine-tuning. Meanwhile, tailoring the responses to the model has been a crucial ingredient for the success of later phases of LLM development, particularly through on-policy preference learning (Tajwar et al., 2024; Zhang et al., 2024c, a; Miao et al., 2024; Gulcehre et al., 2023; Azar et al., 2023a; Tang et al., 2024a; Zhuang et al., 2023), and on-policy/online reinforcement learning (Guo et al., 2024b; Liu et al., 2024d; Zhou et al., 2024c).
Inspired by these insights, we hypothesize that SFT can similarly benefit from aligning data with the model, the core idea behind GRAPE. For each instruction, GRAPE gathers and selects response(s) from various sources that are closest to the target model’s pretrained distribution. This is achieved by calculating the probability of each response using the target model and selects the one with the highest probability(§3).After obtaining these in-distribution responses, one can proceed with standard SFT without any modification to the training. Unlike existing datasets that usually contains one-size-fits-all responses for each instruction without customization (Yu et al., 2024; Yuan et al., 2024; Lian et al., 2023b; Teknium, 2023), GRAPE curates model-dependent SFT datasets that better matches the base model’s distribution, better mitigating the risks associated with distribution shift like spurious correlations and catastrophic forgetting, while posing minimum overhead.
To GRAPE’s advantage, many existing datasets share overlapping instructions but contain different high-quality responses, e.g., the instructions in Flan (Longpre et al., 2023), GSM-8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b), and the post-training recipes that re-use SFT instructions for preference learning (Lambert et al., 2024; OLMo et al., 2025). Therefore, GRAPE can directly select, for each model, the best fit(s) among the off-the-shelf responses without producing new responses, which we do in the experiments(§ 5).
We first validate our hypothesis through extensive controlled experiments on a reasoning dataset - UltraInteract-SFT (Yuan et al., 2024) and demonstrate the importance of supervising base models with in-distribution responses(§4). We experimented on 4 popular pretrained LMs from Mistral (MistralAI, 2024c), Llama3.1 (Dubey et al., 2024) and Qwen2.5 (Hui et al., 2024) families. Notably, models fine-tuned with GRAPE-selected responses outperform those trained on a 3$\times$ larger datasets up to $17.3\%$ absolute gain on average performances, and responses from the strongest model under consideration - Llama3.1-405B-Instruct by significant levels. We then experiment in realistic scenarios, collecting responses from post-training data for Tulu3 and Olmo-v2. We demonstrate the effectiveness of GRAPE by outperforming training on all these available data that is 4.5 times the size of data curated by GRAPE by up to 6.1%. Remarkably, GRAPE allows finetuning a Llama3.1-8B base model to exceed the performance of Tulu-8B-SFT using 1/3 data and half number of epochs.
Our results prove GRAPE as a simple but effective and scalable approach to improving supervised fine-tuning by aligning training distribution and that of the pre-trained checkpoint.
2 Background and Motivation
An analogy from Reinforcement Learning and Preference Learning
The investigation of this work into the distribution match between the pre-trained LM and supervised fine-tuning (SFT) data is inspired by recent findings on policy optimization for LM alignment with RL (Ouyang et al., 2022) and preference learning (Rafailov et al., 2023; Ethayarajh et al., 2024). While the importance of matching training data distribution with the policy has been well noted in both traditional RL (Shi et al., 2023; Fujimoto et al., 2018; Kumar et al., 2019; Peng et al., 2019; Wang et al., 2021; Arora & Goyal, 2023; Jiang & Li, 2016; Tang & Abbeel, 2010) and LM settings (Xiong et al., 2024), preference learning algorithms like DPO (Rafailov et al., 2023), IPO (Azar et al., 2023b) and KTO (Ethayarajh et al., 2024) first emerged as off-policy algorithms. However, subsequent research has highlighted the performance gap between on-policy and off-policy training due to distribution shifts  (Xu et al., 2024b; Tang et al., 2024b) and proposed various mitigation strategies  (Zhuang et al., 2023; Zhou et al., 2024c; Zhang et al., 2024a; Xiong et al., 2024; Guo et al., 2024b), showing that training models on data more closely aligned with their policy distribution can significantly improve performance, while failing to do so can yield sub-optimal policies or those that are harder to generalize.
Hypothesis: Supervised Fine-tuning Benefits From Data That Better Matches Base Distribution
The base distribution of pre-trained language models—shaped by extensive training on vast and diverse datasets—is inherently robust and generalizable. Therefore, during supervised fine-tuning phase, the pre-trained distribution should be carefully preserved (Kumar et al., 2022; Cohen-Wang et al., 2024; He et al., 2023; Yang et al., 2024d; Ding et al., 2023), to best retain the knowledge and capabilities that emerge during pre-training (Zhou et al., 2023). If the proximity between the pre-trained distribution and the fine-tuning data is not maintained, the limited number of training examples available during SFT, compared to the vast scale of pre-training data, can increase the risk of distribution distortion. This misalignment can lead to issues such as catastrophic forgetting (Aghajanyan et al., 2020; Yang et al., 2024d) and the emergence of spurious correlations (Feldman, 2021).
The central premise of our work is that by using responses closely aligned to the pre-trained distribution, we can minimize distribution shift during SFT and therefore achieve better data efficiency and stronger performance.
From On-Policy Alignment To Distribution-Aligned SFT
We build on principles of on-policy alignment techniques with key distinctions tailored for SFT. Given that SFT reprsents a more primitive stage of post-training than preference learning, strictly sampling responses from the base model itself can introduce risks such as instability, bias reinforcement, knowledge stagnation, and overfitting (Herel & Mikolov, 2024; Mobahi et al., 2020; Allen-Zhu & Li, 2023). To address this, we advocate for an approach that stays in-distribution while delivering effective supervision to the base model. To this end, we propose to gather and select responses from various sources, and select one that is closest to the target model’s pretrained distribution, which we name GRAPE.
3 Methodology
Figure 1: An overview of GRAPE. GRAPE takes multiple off-the-shelf existing datasets and optionally generates new responses, finds overlapping instructions with multiple different responses and selects responses that align with the base model’s distribution. This dataset is then used for standard supervised fine-tuning.
We introduce GRAPE, a surprisingly simple yet effective methodology to enhance supervised fine-tuning (SFT) by customizing the training data for the base model. The key idea is to find responses among a candidate pool for each instruction $x_{i}$ such that they align closely with the base model’s pretrained distribution $\pi_{\theta_{0}}$.
As diagrammed in Figure 3, GRAPE consists of two main steps, followed by standard SFT:
  1. 1.
    Response Collection (§3.1) Collect a pool of high-quality candidate responses either from existing datasets or sampling from multiple LLMs.
  2. 2.
    Customization (§3.2): For the target model to be finetuned $\pi_{\theta_{0}}$, find the response(s), for each instruction, that are closest to the pretrained distribution of $\pi_{\theta_{0}}$.
3.1 Collecting Responses from Existing Resources
For instruction-tuning of language models, high-quality instructions are more difficult to collect than responses (Xu et al., 2024c; Liu et al., 2024a). Therefore, it is a common practice to reuse existing instruction-tuning prompts while generating diverse responses using various methods tailored to specific requirements. For instance, instructions from Flan (Longpre et al., 2023), OpenOrca (Lian et al., 2023a), ShareGPT (Team, 2023), and the training splits of GSM-8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b), and CodeContests (Li et al., 2022) are frequently reused in datasets like Olmo (OLMo et al., 2025), Tulu (Lambert et al., 2024), OpenHermes (Teknium, 2023), OpenOrca (Lian et al., 2023a), MetaMath (Yu et al., 2024), MathInstruct (Yue et al., 2023), UltraFeedback (Cui et al., 2024), and UltraInteract (Yuan et al., 2024), whether for SFT or preference learning. The solutions are generated using different models or follow varying styles depending on the specific needs. This naturally leads to a situation where a single instruction with multiple responses becomes a readily available resource. GRAPE therefore leverages these pre-existing response candidates to tailor training dataset that better aligns with the base model’s distribution. When such resources are unavailable or insufficient, practitioners can generate new responses and apply GRAPE.
Responses for the $i$-th instruction are collected from various datasets to form $\mathcal{R}_{i}=\{y{{}_{i}^{j}}:j=1,\dots,J^{(i)}\}$, where $J^{(i)}$ denotes the number of responses collected for the $i$-th instruction.
3.2 Customize Dataset For Models
We then compute the conditional probability of each response $\pi_{\theta_{0}}(y{{}_{i}^{j}}\mid x_{i})$. We rank the responses based on the conditional probability and take those with the highest probability for each instruction.
Notably, GRAPE’s selection process requires only a forward pass through the candidates and does not require gradient computation. Its overhead is substantially lower than many model-based data selection algorithms (Xia et al., 2024; Yang et al., 2024b; Liu et al., 2024b; Zhao et al., 2021; Zhang et al., 2024b; Pan et al., 2024).
Additionally, it is important to distinguish GRAPE from the other preplexity-based data selection and curriculum planning methods (Wu et al., 2024; Li et al., 2024b; Liu et al., 2024c). Existing approaches focus on selecting instructions by using perplexity as a difficulty measure, which differs from GRAPE that uses probability to select for each instruction in a fixed instruction set, responses that better matches with the base model’s distribution. Our experiments in §5 demonstrate that low-probability responses with fixed set of instructions are detrimental to performance, further emphasizing the fundamental difference in the two processes.
4 Preliminary Experiments
In this section, we seek to empirically verify the hypothesis in §2. To evaluate it under a controlled set-up, we consider using instructions from UltraInteract-SFT, and constrain within chain-of-thought reasoning (predominantly for coding, logic and math.) scenario. This allows us to control confounding factors like style and format while ensuring manageable data quality control during data collection. The insights from this experiment will inform the later more realistic setting in §5.
4.1 Experimental Setup
We detail our experiment settings below.
4.1.1 Training Data Curation
In this controlled experiment, we focus on chain-of-thought reasoning (Wei et al., 2022; Wang et al., 2024; Luo et al., 2024; Cobbe et al., 2021; Li et al., 2023; Lightman et al., 2023). Different models may follow different reasoning paths to solve a problem, while their final solutions can be easily verified.
Figure 2: The training loss curve of Llama3.1-8B on Tulu-3-Olmo-2 combo. Best stands for training with highest-probability responses (GRAPE-selected); Random stands for randomly chosen responses for each instruction in the pool, Worst stands for lowest probability responses in the base model’s distribution. Throughout training, Best has lower loss than Random than Worst.
We use UltraInteract-SFT (Yuan et al., 2024), which contains approximately $80,800$ unique instructions covering coding, math (chain-of-thought and program-aided) and logic reasoning domains , where each instruction is paired with varying numbers of different responses to contain a total $>280,000$ training examples. The responses in the dataset are strictly in step-wise format. For each instruction, we use GRAPE to select the top-ranking responses, ensuring that the number of responses matches the original UltraInteract-SFT dataset for fair comparisons.
Response Generation
We collect responses from a diverse set of models of various sizes across model families, including Mixtral-7x7B-Instruct (Jiang et al., 2024), Codestral-22B (MistralAI, 2024a), Mistral-Small (MistralAI, 2024b), Llama-3.1-70B-Instruct and Llama-3.1-405B-Instruct (Dubey et al., 2024), and Qwen2.5-72B-Chat (Yang et al., 2024a), resulting in approximately 10x additional responses per instruction. The responses are then filtered based on the answers to ensure their validity following Yuan et al. (2024).
Figure 3: The average performance curve of Llama3.1-8B on Tulu-3-Olmo-2 combo. Best stands for training with highest-likelihood responses (GRAPE-selected); Random stands for randomly chosen responses for each instruction in the pool. The average is taken over all items on table 2 except AlpacaEval2.
4.1.2 Base Models
To demonstrate the generalizability of GRAPE, we evaluate its performance across multiple LLMs, including Llama-3.1-8B and Llama-3.2-3B from Llama-3 (Grattafiori et al., 2024) family, Mistral-7B (Jiang et al., 2023) and Qwen2.5-7B (Hui et al., 2024).
4.1.3 Evaluation
We evaluate the model on coding and math reasoning benchmarks. For coding tasks, we consider HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), LeetCode (Guo et al., 2024a); for math datasets, we consider MATH dataset (Hendrycks et al., 2021b), GSM-Plus (Li et al., 2024d) and TheoremQA (Chen et al., 2023b) dataset. HumanEval and MBPP are natural-language-to-code benchmarks testing language models’ ability to produce functionally correct programs. LeetCode contains interview-level programming problems that are more challenging. MATH contains high-school level math competition problems, whereas GSM-Plus is a more challenging variant of GSM-8k (Cobbe et al., 2021) and Theorem-QA contains complex math reasoning problems.
4.1.4 Baselines
We compare GRAPE with several baselines. Except for Up-scaled Dataset which is 3×3$\times$3 × larger, the rest are controlled for the number of responses per-instruction to be the same as UltraInteract-SFT.
  • Original dataset. Performing SFT over the original UltraInteract-SFT dataset as is. The original dataset also contains high-quality, verified responses.
    We compare GRAPE to directly using a standard, well-curated dataset. Responses from a powerful model are likely to represent an upper-bound quality of training data, and comparing against this baseline allows us to isolate whether further improvements stem from the strategic response selection performed by GRAPE.
  • Responses from the strongest model under consideration. We use responses exclusively from the best model used to generate responses (in our case, Llama3.1-405B-Instruct).
  • Up-scaled Dataset. The comparison with a larger dataset assesses how well selection improves instruction tuning compared to training on a larger set of diverse correct responses. Each instruction is paired with three times of randomly selected responses from the pool. [^1] For example, if UltraInteract contains 3 different responses for instruction $x$, we end up including 9 for this setting.
We train all models for 1 epoch with a learning rate of $10^{-5}$.
MATH GSMPlus TheoremQA
Model Data HE LC MBPP CoT PoT CoT PoT CoT PoT Avg. Abs. $\Delta$
Original-UI 46.3 15.6 50.1 21.6 32.6 45.9 45.3 16.8 20.1 32.7 3.6
Llama3.1-405B 44.5 12.2 46.9 24.5 33.8 48.0 50.0 17.5 16.8 32.7 3.6
3x Data 48.1 13.3 52.8 25.1 26.1 50.0 49.4 16.8 12.4 32.7 3.6
Mistral-7B Ours 52.4 15.6 53.4 28.9 34.6 50.5 52.8 17.8 20.6 36.3 -
Original-UI 54.3 11.1 58.9 29.7 31.0 53.7 51.6 20.0 20.8 36.8 4.7
Llama3.1-405B 56.7 15.0 60.0 34.8 38.1 51.4 55.4 16.6 21.0 38.8 2.7
3x Data 48.8 7.8 57.9 25.5 11.2 48.6 45.1 20.6 19.6 31.7 9.8
Llama3.1-8B Ours 57.3 19.4 63.8 34.8 39.2 56.6 56.1 22.5 23.9 41.5 -
Original-UI 32.9 3.9 41.6 12.8 16.1 30.8 19.5 14.6 10.5 20.3 3.8
Llama3.1-405B 31.7 5.0 43.3 6.6 6.6 30.8 20.6 15.1 10.8 18.9 5.1
3x Data 42.6 6.7 42.9 8.7 5.1 17.8 19.5 14.6 12.8 19.0 5.1
Llama3.2-3B Ours 42.6 13.3 44.6 16.4 17.6 34.9 20.6 15.1 11.4 24.1 -
Original-UI 67.0 41.2 60.0 51.0 38.3 64.1 59.5 14.6 10.5 45.1 11.6
Llama3.1-405B 71.3 45.0 62.0 31.5 35.1 40.1 36.4 33.3 31.2 42.9 13.8
3x Data 75.6 48.3 62.9 32.8 24.5 47.1 22.8 21.6 19.0 39.4 17.3
Qwen2.5-7B Ours 77.4 48.9 70.7 56.4 45.3 67.7 66.3 37.4 40.1 56.7 -
Table 1: Result of synthetic experiment on UltraInteract-SFT. The last column, Abs. $\Delta$ is the absolute improvement of GRAPE over that row.
4.2 Results and Analysis
Table 1 summarizes the performance of GRAPE across benchmarks. Our approach consistently outperforms the various baselines across the board, including the original UltraInteract-SFT dataset. GRAPE-selected solutions can outperform those directly sampled from the strongest model under consideration (Llama3.1-405B-Instruct) up to 13.8% absolute improvement. This implies that customization for base models should be prioritized over identifying the presumably highest-quality responses. This verifies our central premise that being in-distribution with each base model is an important ingredient for the responses we supervise the base models on, to boosting downstream performance. Furthermore, we demonstrate that merely adding more responses does not always lead to continuous improvement in model performance, which aligns with findings in prior studies (Li et al., 2024c; Du et al., 2023). By properly aligning with models’ base distributions, GRAPE outperforms those trained with 3x responses with at least 3.6% and up to 17.3% absolute improvement. These results reinforce the notion that scaling data without considering its alignment with the base model’s initial distribution risks diminishing returns and, in some cases, even performance degradation.
5 GRAPE-Picking From Real-World SFT Datasets
In this section, we leverage the findings from the earlier experiments and demonstrate the effectiveness of GRAPE to customize training data for each base model by selecting from available datasets with overlapping instructions. Here, we do not generate any new responses for the instructions; it only selects from existing ones. We evaluate GRAPE on the fully open dataset used in post-training phases of Tulu-3 (Lambert et al., 2024) and Olmo-2 (OLMo et al., 2025). The details are presented below.
5.1 Data Mixture Details
Tulu-3 (Lambert et al., 2024) is a fully open-source collection of post-training recipes, including supervised fine-tuning and preference alignment data. Olmo-2 (OLMo et al., 2025) is a fully open-source language model. Both Tulu-3 and Olmo-2 use the same data mixture during the supervised fine-tuning stage, but different data mixtures and source models for generating preference data for different sizes of their models: Tulu-3-8B/70B and Olmo-2-7B/13B. To demonstrate the effectiveness of GRAPE, we collected the overlapping instructions from both models and gather their corresponding responses. From the preference data, we retained only the winning responses. We formed our candidate pool with those instructions with at least two distinct responses, resulting in a dataset of 350.4k unique instructions and about 1.03 million total instruction-response pairs for evaluation with GRAPE. We do not apply further processing of these data or any filtering on top of GRAPE.
5.2 Evaluation
We evaluate on a set of commonly used benchmarks spanning over coding, math, knowledge and instruction-following. We evaluated on LeetCode (Guo et al., 2024a), MATH (Hendrycks et al., 2021b), BigBenchHard(BBH) (Suzgun et al., 2022), MMLU (Hendrycks et al., 2021a), and AlpacaEval-V2 (Dubois et al., 2024). LeetCode, MATH, BBH and MMLU are evaluated the in the same way as in (Yuan et al., 2024), where we use zero-shot for MATH and MMLU, 3-shot example for BBH. We use the same AlpacaEval-v2 as in OpenInstruct.
5.3 Baselines
  • Responses From The Original SFT on The Subset. In this experiment, we keep the same set of instructions (350.4k) and pair each instruction with the response from the SFT mixture.
  • Random Response Candidate From The Pool. Randomly pairing each instruction with a candidate response for that instruction, which may come from either the SFT mixture or a winning response in the preference-learning data.
  • Lowest-Probability Response. Instead of taking the highest probability response to each instruction for the base model, we do the opposite on and take the lowest for each.
  • Entire SFT Dataset. We train on the entire SFT data mixture of Tulu with 939k instances.
  • All Available Responses For The Overlapping Instructions. In this setting, we train on the entire candidate pool for all overlapping instructions, resulting in 103.6k instances.
  • S2L (Yang et al., 2024b). S2L is a state-of-the-art unsupervised data selection baseline designed to select balanced subsets from large datasets by leveraging training loss trajectories. It first trains a small reference model within the same model family and record loss trajectory for each training example, then applies K-means clustering on the loss trajectory and samples equally from each. We use Llama-3.2-1B as reference for Llama3.1-8B and Qwen2.5-0.5B for Qwen2.5-7B. For Mistral-v0.3-7B, we use itself as reference since no smaller models in the family are available. For S2L, we select the same number of data from the same pool as GRAPE. Further details in Appendix B.
  • All data available for SFT mixture. We train with all the available responses from all these datasets that share prompts with the SFT mixture. Which gives approximately 1.58 million (4.5 times thesize of GRAPE-curated dataset). This is the union of the entire dataset and GRAPE’s candidate pool.
AlpacaEval2
Model Data Num. Instances LC WR BBH MMLU MATH LeetCode Avg. Abs. $\Delta$
Highest 350.4k 10.1 6.2 68.9 63.2 22.9 13.3 30.8 5.2
Random 350.4k 12.8 10.8 68.6 63.1 27.9 13.3 32.8 3.2
SFT-Only 350.4k 7.1 5.5 68.9 64.1 20.2 17.2 30.5 5.4
Tulu3-SFT 939k 12.4 8.0 67.9 65.9 31.5* 7.8 32.4 3.5
All Responses 1.03M 12.9 11.4 68.7 62.8 32.1 17.2 34.2 1.8
S2L 350.4k 8.5 7.6 68.6 63.1 26.5 16.1 31.7 4.2
All Available Data 1.58M 8.8 10.1 69.8 62.1 32.5 16.1 33.2 2.7
Llama 3.1-8B GRAPE 350.4k 14.8 15.2 69.6 64.5 32.1 19.4 35.9 -
Highest 350.4k 7.0 5.5 58.8 56.1 15.1 10.6 25.5 6.4
Random 350.4k 10.6 9.2 60.0 57.8 19.6 11.1 28.1 3.9
SFT-Only 350.4k 7.1 5.3 53.9 57.0 14.4 12.0 24.9 7.0
Full-SFT-Data 939k 11.5 10.5 59.0 55.2 25.8 15.6 29.9 2.0
All Responses 1.03M 10.5 11.5 61.0 57.9 24.2 14.4 29.9 2.0
S2L 350.4k 10.5 11.9 61.9 57.1 22.4 13.9 29.6 2.3
All Available Data 1.58M 8.0 7.0 55.3 53.7 25.4 12.3 26.9 5.0
Mistral -7B GRAPE 350.4k 13.6 13.9 62.3 59.2 24.2 18.3 31.9 -
Highest 350.4k 8.0 10.7 72.2 73.2 49.4 42.2 42.6 5.9
Random 350.4k 16.1 14.9 73.3 73.1 56.0 43.3 46.1 2.4
SFT-Only 350.4k 9.9 7.8 71.2 74.1 51.1 46.6 43.4 5.1
Full-SFT-Data 939k 9.5 7.1 71.4 73.1 47.0 48.3 42.7 5.8
All Responses 1.03M 16.0 14.5 71.4 72.1 51.7 43.3 44.8 3.7
S2L 350.4k 13.4 14.9 72.7 73.1 53.4 40.6 44.7 3.9
All Available Data 1.58M 13.3 12.3 70.3 71.8 44.0 42.8 42.4 6.1
Qwen2.5 -7B GRAPE 350.4k 20.0 20.4 73.2 73.3 60.0 44.4 48.6 -
Table 2: GRAPE on the Tulu-Olmo collection. For Llama3.1-8B base model, we included Tulu3-SFT model’s results. The “*”-marked number for MATH is obtained using 4-shot prompting. We train all the models (except that we took Tulu3-SFT numbers directly) for 1 epoch with a learning rate of $10^{-5}$.
5.4 Results
As shown in Table 2, models fine-tuned on responses selected by GRAPE outperforms the strong baselines we constructed, especially the one that trains over all available data by significant margins across the 3 models. Remarkably, Using roughly 1/6 training computation (Tulu3-8B-SFT was trained for 2 epochs on 3 times of data), our performance exceeds that of Tulu3-8B-SFT. Also, GRAPE outperforms state-of-the-art data-selection approaches like S2L, despite its simplicity and efficiency, further highlighting its effectiveness in diverse real-world scenarios and strengthening its overall practicability for large-scale instruction tuning and robust generalization. Without the need to synthesize any new data, one can easily leverage established datasets sourced from the web to customize a dataset for each base model that yields better fine-tuning outcome.
These results feature GRAPE not only as an effective strategy to enhance performances, but a handy approach to improve fine-tuning efficiency.
5.5 Further Experiments
Item
Metric Data
Llama
3.1-8B
Mistral
-7B-v0.3
Qwen
2.5-7B
Subset 8.6 5.9 7.6
Random 8.0 6.2 9.0
LC GRAPE 11.3 8.2 10.8
Subset 6.2 3.9 5.2
Random 6.4 4.8 7.2
Alpaca -Eval2 WR GRAPE 9.4 7.5 9.6
Subset 51.6 49.0 54.4
Random 51.4 49.9 55.2
Truthful -QA MC2 GRAPE 52.7 51.6 56.4
Table 3: Results on OpenHermes-2.5. The Subset row refers to training exclusively on the SFT responses over the subset.
We extend our evaluation using OpenHermes-2.5 (Teknium, 2023), a high-quality, open instruction-tuning dataset containing approximately 1 million distinct instructions. Following a similar strategy to the previous combo experiment, we source responses from additional datasets from  Huang et al. (2024) and HuggingFace-H4 (2024). For preference-based datasets, we include only the winning responses, consistent with the methodology in the Tulu-Olmo experiment. This process results in 575K unique instructions and 1.34 million total instruction-response pairs. As shown in Table 3, GRAPE-selected data yields better performances too. The consistent improvement across models reaffirm our earlier observations: GRAPE serves as an adaptive response selection mechanism that significantly enhances SFT performance that has wide applicability.
5.6 Discussion:GRAPE works if all responses are from the same LM
GRAPE
Qwen2.5
-72B
Llama3.1
-405B
Gemma
-it-9B
LC 28.1 25.8 16.3 26.9
WR 33.1 24.3 17.0 20.2
Table 4: Alpaca-Eval2 Results On Magpie-Zoo (Xu et al., 2024d).
In Section 4, we demonstrated how GRAPE enables practitioners to refine model-generated responses for improved training outcomes. Beyond that, GRAPE can optimize responses from a single generator. Using the Magpie-Zoo instruction set (Xu et al., 2024d), we sample 10 responses per instruction from Qwen2.5-72B-Instruct, select in-distribution responses with GRAPE, to train a Mistral-v0.3-7B model. We also compare with training on top-3 reported datasets of Magpie-Zoo. As shown in Table 4, GRAPE-selected responses further boost the performance. Practically, batch sampling of responses cause little latency (Zhong et al., 2024; Zhou et al., 2024d). One therefore can sample batches of responses from a single strong model and apply GRAPE on top of it to find the response closest to the base model’s distribution efficiently.
5.7 Discussion: Is In-Distribution A Silver Bullet?
The success of GRAPE verifies our key hypothesis that matching SFT distribution to the base model benefits the performance. Yet, we argue that this cannot be pushed to the limit of SFT using data from the same distribution. By experimenting with training on solutions sampled from a trained version of the same base model, we notice a drastic performance drop from the original model as shown in Table 5.
Model Data Avg.
Self-Distilled 28.4(-)
Mistral-7B Original-UI 32.7
Self-Distilled 29.4(-)
Llama3.1-8B Original-UI 36.8
Self-Distilled 15.1(-)
Llama3.2-3B Original-UI 20.3
Table 5: Performance Degradation From Self-Distillation On UI.
We explored this using the MATH (Hendrycks et al., 2021b) dataset. A base model was fine-tuned using solutions from stronger models like Llama3.1-70B-Instruct or GPT-3.5-Turbo-augmented MetaMathQA (Yu et al., 2024). This fine-tuned model then generated new solutions, which were used to further fine-tune another base model instance, simulating iterative on-policy tuning.
Dataset Model Response Generator N Acc.
MATH Mistral Llama3.1-70B 10 18.2
MATH Mistral FT-Mistral 10 15.9 (-)
MATH Llemma Llama3.1-70B 10 26.2
MATH Llemma FT-Llemma 10 23.6 (-)
MATH Mistral MM-AnsAug - 22.3
MATH Mistral FT-Mistral 10 20.6 (-)
MATH Llemma MM-AnsAug - 28.1
MATH Llemma FT-Llemma 10 21.4 (-)
Table 6: Effect of self-generation on MATH dataset. FT-Mistral refers to the model right above that row finetuned from either MM-AnsAug or Llama3.1-70B-Instruct produced solutions. N stands for the number of responses sampled.
As shown in Table 6, performance declined when solutions were sampled from the fine-tuned model itself. This decline stems from reduced solution diversity, leading to distributional drift and poorer generalization.
These results validate our claim in §2 that robust fine-tuning requires more than alignment with the base model—it needs complementary strategies to maintain response diversity and quality.
6 Related Works
Data Engineering For Instruction Tuning
Data is central to the success of effective instruction tuning,  (Xu et al., 2023; Xia et al., 2024; Chan et al., 2024), featuring both automated data synthesis (Xu et al., 2024a; Zeng et al., 2024; Yu et al., 2024; Wei et al., 2023) and selection (Xia et al., 2024; Chen et al., 2023a; Parkar et al., 2024; Li et al., 2024e). Some selection approaches focus on high-quality data by leveraging LLMs (Chen et al., 2023a; Parkar et al., 2024; Li et al., 2024b) or employing principled metrics (Kang et al., 2024; Mekala et al., 2024; Xia et al., 2024), while others, such as  Yang et al. (2024c); Das & Khetan (2023), aim to identify diversity-optimized subsets for greater efficiency.
Another emerging trend is the customization of training data based on the characteristics of the base models. For instance,  Li et al. (2024c); Du et al. (2023) leverage the base model itself to select a subset of instructions, while  Li et al. (2024a) introduce a teacher model to guide the selection process. However, these approaches primarily focus on identifying a re-weighted subset of questions for training. In contrast, GRAPE has a different focus, where it aims to find a set of responses from the solution space that not only provide good coverage of the golden distribution but also align closely with the base model’s policy.
On-Policy Methods For Language Model Alignment
Recent advances in RLHF (Ouyang et al., 2022), RLAIF (Lee et al., 2024b) and preference learning (Rafailov et al., 2023) research have identified the benefits of on-policy data for model alignment (Xu et al., 2024b; Tajwar et al., 2024; Zhang et al., 2024c; Guo et al., 2024b; Liu et al., 2024d; Zhang et al., 2024a; Miao et al., 2024; Gulcehre et al., 2023). However, these works primarily emphasize the preference-tuning stage, where models are refined to generate samples from an already reasonable policy. In contrast, less attention is given to the supervised fine-tuning phase, which starts with base models that are not yet well-prepared to generate high-quality answers that align sufficiently with the desired policy for optimization.
7 Conclusion
We present GRAPE, a simple yet highly effective approach to improve supervised fine-tuning data building on our empirically-verified hypothesis that instruction tuning data shall better match the base model’s distribution to optimize the training outcome. GRAPE is surprisingly simple and efficient, where it customizes response data to supervise each model by taking highest-probability instances. We demonstrated the effectiveness of GRAPE on multiple settings. Remarkably, GRAPE can outperform not only SOTA data selection baselines, but also Tulu3-SFT by using a 1/3 subset of its SFT instructions, and performs better than training on 4.5 times data it selects from. We show GRAPE as a highly scalable and simple approach can be adopted in multiple practical scenarios with minimal overhead.
Appendix A Parameter Distance
We measure the L2 norm of model parameter difference between fine-tuned and pre-trained checkpoints, as a signal of how much the distribution has drifted during SFT (ichi Amari, 2016; Cover & Thomas, 2006). We notice that training over well-matched distribution shifts the parameter less than training over those ill-matched.
Mistral-7B-v0.3 Llama3.1-8B Qwen2.5
GRAPE 8.006 8.196 8.426
Worst 8.029 8.202 8.467
Table 7: Performance comparison across different models
Appendix B Further Details On S2L
This section details the experimental setup for two unsupervised data selection baseline, S2L applied to the joined pool of Tulu-3 and Olmo-v2 data.
S2L, a state-of-the-art unsupervised data selection baseline, operates through two key steps: training a reference model to capture training dynamics and clustering the resulting trajectories to form a diverse, balanced subset of training data. The reference models used in our setup are specifically selected to enhance S2L’s performance, adhering to the theoretical underpinnings from the original paper that training dynamics remain consistent across models of varying sizes within the same family.
For our experiments, we train small reference models corresponding to the final target models. Specifically, we pair Llama-3.1-8B with Llama-3.2-1B, Qwen-2.5-7B with Qwen-2.5-0.5B, and Mistral-v0.3-7B with itself due to the lack of smaller models in the Mistral family. To minimize computational costs, LoRA is applied when training the Mistral reference model. This choice of reference models are better compared to original S2L setup, which employed a Pythia-70M proxy, thereby improving the fidelity of the selected subset.
Following S2L, the reference models are trained on a random 5% subset of the dataset over four epochs. This reduced training requirement is justified by prior work, which demonstrates that only partial data is sufficient for the proxy model to learn meaningful training dynamics. During trajectory collection, we record the training loss of all examples at intervals of 500 iterations. The batch size and learning rate schedules are set as batch size of 128 and a learning rate warmup of 3%, followed by a cosine decay to 2e-5.
We then perform K-means clustering using the Faiss library to efficiently partition the trajectory space into 100 clusters. The number of iterations is set to 20, and we use the Euclidean distance metric to ensure convergence to well-separated clusters. From each cluster, an equal number of examples are sampled to maintain a balanced subset distribution.
Appendix C Further Training Details
We train our models on a 4-GPU Nvidia-GH200 node, with batch size 256 and micro batch size 2.
Appendix D Further Ablations on UltraInteract.
Data Full UI Closest-1 Random-1
Num. Instances 280K  80K  80K
HumanEval 46.3 42.1 41.5 (-)
LeetCode 15.6 13.9 11.1 (-)
MBPP 50.1 52.1 49.1 (-)
CoT 21.6 19.2 15.5 (-)
MATH PoT 32.6 24.9 15.1 (-)
CoT 45.9 44.1 35.3 (-)
GSMPlus CoT 45.3 43.2 45.2 (-)
CoT 16.8 15.8 15.8
TheoremQA PoT 20.1 12.9 15.3
Avg. 32.7 29.8 27.1(-)
Table 8: Ablations on data selection with Mistral-7B-V0.3 by selecting within UltraInteract-SFT (since it contains varying numbers of responses per-instruction). Closest-1 denotes the one closest to the base model’s initial distribution. Random-1 is sampled from the entire enlarged dataset formed by both original and generated responses. We use (-) to denote Random-1 underperforming Closest-1.
Model Data HE LC MBPP MATH GSMPlus TheoremQA Avg.
CoT PoT CoT PoT CoT PoT
Mistral-7B Self-Distill 46.3 13.3 49.6 17.3 18.5 43.3 33.2 16.8 17.4 28.4
Original-UI 46.3 15.6 50.1 21.6 32.6 45.9 45.3 16.8 20.1 32.7
Ours 52.4 15.6 53.4 28.9 34.6 50.5 52.8 17.8 20.6 36.3
Llama3.1-8B Self-Distilled 47.6 6.7 51.7 22.9 12.7 47.2 35.3 18.8 21.5 29.4
Original-UI 54.3 11.1 58.9 29.7 31.0 53.7 51.6 20.0 20.8 36.8
Ours 57.3 19.4 63.8 34.8 39.2 56.6 56.1 22.5 23.9 41.5
Llama3.2-3B Self-Distilled 32.3 5.6 41.9 8.8 7.0 12.1 12.1 5.9 10.5 15.1
Original-UI 32.9 3.9 41.6 12.8 16.1 30.8 19.5 14.6 10.5 20.3
Ours 42.6 13.3 44.6 16.4 17.6 34.9 20.6 15.1 11.4 24.1
Table 9: The detailed comparison across benchmarks for self-distillation discussed in Section 5.7
Appendix E Additional Related Works On Model Dependent Data Selection Approaches
E.1 0. Notations
  1. 1.
    A training dataset $D=\{x_{i}\}_{i=1}^{N}$ of size $N$; the final language model to be trained on the selected data $\theta$.
  2. 2.
    We denote the average cost of one forward pass of model $\theta$ on a training example as $F_{\theta}$. As one backward pass is approximately the cost of two forward passes, the average cost of one “gradient pass” (i.e., one forward + one backward) is thus $3F_{\theta}$.
  3. 3.
    Another important source of computational cost in data selection comes from the training of additional models. We use $C(\theta,D,T)$ to denote the cost of training model $\theta$ on dataset $D$ for $T$ epochs (i.e., $N\cdot T$ examples are seen in total).
  4. 4.
    Therefore, we unify the computational cost of most data selection approaches into two parts:
    1. (a)
      The training of additional models. For example, gradient-based influence requires training an additional model on part of the training dataset for $T$ epochs to obtain the checkpoints for gradient computation.
    2. (b)
      The computation of per-sample features. For example, for each training example, gradient-based influence requires computing its gradient for each saved checkpoint, which means $T$ gradient passes are needed.
  5. 5.
    Note that some algorithms may have additional computational costs other than the two parts above, such as clustering or a greedy algorithm for the final data selection. Since the two parts above constitute the majority of computation for almost all the data selection approaches, we omit the other cost and only focus on these two.
1. TLDR: The Final Table
For GRAPE, we assume that in the training dataset $D$, various responses to the same instruction are already available, thus no additional cost is incurred in the Response Collection step of GRAPE. So the computational cost analysis of GRAPE under our framework is:
  • Additional Training: 0, as GRAPE directly evaluates data using the base model.
  • Per-sample conditional probability: $NF_{\theta}$, as for a given target model $\theta$, we only need to compute conditional probability for each response (example) once.
The table below shows that our method, GRAPE, achieves superior performance with minimal computational cost compared with other model-based data selection approaches.
Additional Training Per-Sample Feature Computation
GRAPE (ours) 00 $NF_{\theta}$
Gradient-based influence (LESS) $C(\theta_{\text{lora}},D_{\text{warmup}},T)$ $3T\cdot NF_{\theta}$
In-run gradient-based influence $C(\theta,D,1)$ 00
Gradient matching $C(\theta_{\text{lora}},D_{\text{warmup}},T)$ $3T\cdot NF_{\theta}$
Gradient norm $m\cdot C(\theta,D,1)$ $3m\cdot NF_{\theta}$
Embedding-based 00 $NF_{\theta}$
Simple uncertainty indicators 00 $NF_{\theta}$
Perplexity $C(\theta_{\text{ref}},D_{\text{ref}},1)$ $NF_{\theta_{\text{ref}}}$
Learnability $C(\theta,D,1)$ $2\cdot NF_{\theta}$
Loss trajectory (S2L) $C(\theta_{\text{ref}},D,T)$ $T\cdot NF_{\theta_{\text{ref}}}$
Table 10: Computational cost comparison of data selection methods.
2. Gradient-based Methods
Gradients have long been an important source of information for training data selection, as they directly affect the whole optimization process of language models. Three kinds of model-based gradient-based data selection approaches have been proposed:
  1. 1.
    Gradient-based influence
  2. 2.
    Gradient matching
  3. 3.
    Gradient norm
2.1 Gradient-based Influence
Gradient-based influence computes the pairwise influence scores between each pair of training and validation examples. Training data with the highest influence are selected, as training on them leads to the theoretically largest decrease in model loss on validation data. LESS (Xia et al., 2024) formulates the pairwise influence scores as the cosine similarity between the gradients of training and validation data, and computes these gradient features using the following two steps:
  1. 1.
    LoRA-train the final model on part of the whole training dataset, denoted as $D_{\text{warmup}}$, for $T$ epochs, and save the $T$ model checkpoints.
  2. 2.
    For each data point, compute its LoRA gradient with each of the $T$ checkpoints, and later aggregate these $T$ gradients together in the cosine similarity expression.
Therefore, the computational cost of gradient-based influence is:
  • Additional training: $C(\theta_{\text{lora}},D_{\text{warmup}},T)$.
  • Per-sample gradient for each checkpoint: $NT\cdot 3F_{\theta}=3T\cdot NF_{\theta}$.
In order to reduce the cost incurred by per-sample gradient computation, recent work has developed in-run gradient-based influence that directly computes the dot product between gradients without the need for separate gradient computations. However, this approach incorporates the dot product computations into the standard training process, which means in order to obtain pairwise influence scores for the whole training set, a full training run has to be done on all the training data. This incurs inefficiency when we do not actually need full dataset training. Moreover, the pairwise scores here only show the model’s “dynamic preference”: scores computed at the $t$-th iteration only reflect the model’s preference at this specific iteration. It is not theoretically guaranteed that these scores reflect the model’s preference from the beginning of training. Thus, the cost of in-run gradient-based influence is:
  • Additional Training: $C(\theta,D,1)$.
  • Per-sample gradient: 0.
2.2 Gradient Matching
Gradient matching also requires per-sample gradients, but utilizes their information in a different way. It performs clustering based on these gradient features to group similar data, and then applies an iterative greedy selection algorithm. In order to scale to LLM-level gradient computation and clustering, TAGCOS (TAGCOS) completely follows the warmup training and gradient computation pipeline of LESS (LESS). As the computational bottleneck here is still the gradient computation instead of clustering or iterative selection, (TAGCOS) also shares the same computational cost as (Xia et al., 2024):
  • Additional training: $C(\theta_{\text{lora}},D_{\text{warmup}},T)$.
  • Per-sample gradient for each checkpoint: $NT\cdot 3F_{\theta}=3T\cdot NF_{\theta}$.
2.3 Gradient Norm
The $L_{2}$-norms of gradient vectors can also serve as effective indicators for data selection. (Paul et al., 2023) proposes GraNd, which obtains a utility score for each training point based on its gradient norm early in the training. More specifically, it starts from $m$ different model weight initializations, trains each model on the whole dataset to obtain per-sample gradient norms, and finally averages the $m$ gradient norms for each training point to obtain the final GraNd score. Therefore, the computational cost of GraNd is shown below:
  • Additional training: $m\cdot C(\theta,D,1)$.
  • Per-sample gradient for each weight initialization: $Nm\cdot 3F_{\theta}=3m\cdot NF_{\theta}$.
3. Embedding-based Methods
Embedding-based methods project the whole training set into an embedding space to quantify the information of each data point and their interactions. For model-based embedding-based selection methods, the embeddings are usually computed by the final model $\theta$ to align with its preference.
Under a supervised data selection setup where validation data representing target task distributions are available, Representation-based Data Selection (RDS; (Rubin et al., 2022; Hanawa et al., 2021)) computes the embedding similarity between training and validation data, and selects training points that are most similar to the target distribution in the embedding space.
For an unsupervised setup where only the embeddings of training data are accessible, geometry-based coreset sampling methods are widely used (Qin et al., 2024). Grounded on the intuition that close samples in the embedding space often share similar properties, a diverse subset can be obtained by controlling the minimum distance between any two selected data points. Among them, using K-center greedy sampling to select embedding-based facility locations has been proven especially effective for instruction fine-tuning of LLMs (Bhatt et al., 2024a).
These embedding-based approaches share similar computational costs: they do not need any additional model training and can directly extract useful per-sample embeddings using the last-layer hidden states of the pretrained final model $\theta$. Thus, their computational cost is shown below:
  • Additional training: 0.
  • Per-sample embedding computation: $NF_{\theta}$.
4. LogProb-based Methods
LogProb-based methods also directly utilize the target LLM to evaluate the utility of each training data point.
4.1 Simple Uncertainty-based Indicators
Some simple model-based indicators inspired by the notion of uncertainty have been shown effective for a long time and recently extended to data selection for LLM instruction tuning (Marion et al., 2023b; Bhatt et al., 2024b). (Bhatt et al., 2024b) demonstrates the effectiveness of various indicators including mean entropy, least confidence, mean margin, etc. These simple indicators do not require additional training and can also be directly obtained with the pretrained final model $\theta$. Their computational cost is shown below:
  • Additional training: 0.
  • Per-sample per-token logits computation: $NF_{\theta}$.
4.2 Perplexity (PPL)
PPL is also a long-standing data selector and has been shown effective for LLM-scale data selection. Typically, a split of the training dataset, $D_{\text{ref}}$, is needed to train $\theta_{\text{ref}}$, a reference model that will be used to compute PPL for the whole training set.
A common approach is to use the final model $\theta$ as the reference model $\theta_{\text{ref}}$ to ensure the alignment in PPL patterns (Marion et al., 2023a), but prior work (Ankner et al., 2024) also shows that a reference model much smaller than the final model can also be an effective PPL-based data selector. The computational cost for PPL-based selection is shown below:
  • Additional training: $C(\theta_{\text{ref}},D_{\text{ref}},1)$.
  • Per-sample PPL computation: $NF_{\theta_{\text{ref}}}$.
4.3 Learnability
In addition, learnability (Mindermann et al., 2022; Zhou et al., 2024a, b) is a more effective metric than pure uncertainty or PPL, as it excludes uncertain but unlearnable points (e.g., noisy or less task-relevant) by considering the decrease in per-sample loss before and after the model is fully trained. More specifically, it trains the final model $\theta$ on the full training dataset to obtain a strong reference model $\theta_{\text{ref}}$, and then computes the difference of loss on each training example between $\theta$ and $\theta_{\text{ref}}$. In this way, it requires two forward passes for per-sample computation:
  • Additional training: $C(\theta,D,1)$.
  • Per-sample learnability computation: $2\cdot NF_{\theta}$.
4.4 Loss Trajectory
Moreover, logprob-based methods can also obtain finer-grained information from the training dynamics of LLMs. S2L (Yang et al., 2024b) obtains a feature vector for each training point by collecting their training loss trajectories over $T$-epoch training on a small reference model $\theta_{\text{ref}}$, and then applies K-means clustering to equally sample data points from each trajectory cluster. Prior work shows its superiority over other logprob-based indicators, but it also comes with significant computational cost:
  • Additional training: $C(\theta_{\text{ref}},D,T)$. Here the choice of $\theta_{\text{ref}}$ is especially important, as prior work (Xia et al., 2023) shows that reference models that come from the same model family as the final model tend to have similar loss trajectories of training data, so they can preserve more fidelity in their loss trajectory patterns.
  • Per-sample loss trajectory computation: $T\cdot NF_{\theta_{\text{ref}}}$. Note that $T$ here is typically much larger than that in gradient-based influence computation, so the computational cost of this gradient-free approach can be even higher than gradient-based methods.

### v3

@@
 Abstract
-High-quality supervised finetuning (SFT) data are crucial for eliciting strong capabilities from pretrained large language models (LLMs).
-Typically, instructions are paired with multiple responses sampled from other LLMs, which are often out of the distribution of the target model to be finetuned.
-This, at scale, can lead to diminishing returns and even hurt the models’ performance and robustness.
-We propose GRAPE, a novel SFT framework that accounts for the unique characteristics of the target model.
-For each instruction, it gathers responses from various LLMs, and selects the one with the highest probability measured by the target model, indicating that it aligns most closely to the target model’s pretrained distribution; it then proceeds with standard SFT training.
-We first evaluate GRAPE with a controlled experiment, where we sample various solutions for each question in UltraInteract from multiple models and finetune commonly used LMs like Llama3.1-8B, Mistral-7B and Qwen2.5-7B on GRAPE-selected data.
-GRAPE significantly outperforms strong baselines, including distilling from the strongest model with absolute gain up to 13.8% averaging across benchmarks, and training on 3$\times$ more data with maximum 17.3% performance improvements.
-GRAPE’s strong performance generalizes to realistic settings.
-We experiment with the post-training data used for Tulu3 and Olmo-2.
-GRAPE can outperform strong baselines with 4.5 times the data by 6.1% and state-of-the-art data selection approach by 3.9% on average performance.
-Remarkably, using 1/3 data and half number of epochs, GRAPE allows Llama3.1-8B to surpass the performance of Tulu3-SFT by 3.5%.
-Machine Learning, ICML
+[…]
 1 Introduction
-High-quality, large-scale supervised data is crucial for supervised instruction finetuning (SFT; Databricks, 2023; Köpf et al., 2023; Zhao et al., 2024a; Zheng et al., 2024).
-A common practice of data collection involves sampling responses from strong language models, predominantly focusing on expanding the size of the dataset and improving the overall quality of the responses (Sun et al., 2023; Taori et al., 2023; Wang et al., 2023; Xu et al., 2024c; Chen et al., 2024).
-However, recent research suggests that there is more complex dynamics involved (Xu et al., 2024d).
+High-quality, large-scale supervised data is crucial for supervised fine-tuing(SFT; Databricks, 2023; Köpf et al., 2023; Zhao et al., 2024a; Zheng et al., 2024).
+A common practice of collecting SFT data involves sampling responses from strong language models, predominantly focusing on expanding the size of the dataset and improving the overall quality of the responses (Sun et al., 2023; Taori et al., 2023; Wang et al., 2023; Xu et al., 2024b; Chen et al., 2024).
+However, recent research suggests that there is more complex dynamics involved (Xu et al., 2024c).
 A plateau effect in synthetic data scaling, where performance either stagnates or even declines as the size of the synthetic data increases beyond a certain point, has been widely observed.
 This phenomenon arises due to issues such as diminishing diversity (Padmakumar & He, 2024; Guo et al., 2023) and distortion in the data distribution (LeBrun et al., 2021), which ultimately undermine the base model’s performance and robustness (Alemohammad et al., 2024; Gerstgrasser et al., 2024; Shumailov et al., 2023; Dohmatob et al., 2024; Hataya et al., 2023; Martínez et al., 2023a, b; Bohacek & Farid, 2023; Briesch et al., 2023).
-Thus, effective instruction tuning requires more than scaling up the data; it often needs “tailoring” the data to the unique characteristics of the target model.
-Existing works focus on enhancing the model’s existing knowledge and capabilities (Du et al., 2023) and optimizing the curriculum progression for instruction tuning (Zhao et al., 2024b; Lee et al., 2024a; Feng et al., 2023; Setlur et al., 2024).
-They typically prescribe questions for each model to learn solutions.
-However, it remains elusive which responses best suit a given base model for fine-tuning.
-Meanwhile, tailoring the responses to the model has been a crucial ingredient for the success of later phases of LLM development, particularly through on-policy preference learning (Tajwar et al., 2024; Zhang et al., 2024c, a; Miao et al., 2024; Gulcehre et al., 2023; Azar et al., 2023a; Tang et al., 2024a; Zhuang et al., 2023), and on-policy/online reinforcement learning (Guo et al., 2024b; Liu et al., 2024d; Zhou et al., 2024c).
+Thus, effective SFT requires more than scaling up the data; it often needs “tailoring” the data to the unique characteristics of the target model.
+Existing works focus on enhancing the model’s existing knowledge and capabilities (Du et al., 2023) and optimizing the curriculum progression for instruction tuning (Zhao et al., 2024b; Lee et al., 2024; Feng et al., 2023; Setlur et al., 2024).
+They typically find questions the model should best learn, instead of what answers the model should imitate.
+Meanwhile, tailoring the responses to the target model has been a crucial ingredient for the success of later phases of LLM development, particularly through on-policy preference learning (Tajwar et al., 2024; Zhang et al., 2024c, a; Miao et al., 2024; Gulcehre et al., 2023; Azar et al., 2023; Tang et al., 2024a; Zhuang et al., 2023), and on-policy/online reinforcement learning (Guo et al., 2024b; Liu et al., 2024d; Zhou et al., 2024c).
 Inspired by these insights, we hypothesize that SFT can similarly benefit from aligning data with the model, the core idea behind GRAPE.
 For each instruction, GRAPE gathers and selects response(s) from various sources that are closest to the target model’s pretrained distribution.
-This is achieved by calculating the probability of each response using the target model and selects the one with the highest probability(§3).After obtaining these in-distribution responses, one can proceed with standard SFT without any modification to the training.
-Unlike existing datasets that usually contains one-size-fits-all responses for each instruction without customization (Yu et al., 2024; Yuan et al., 2024; Lian et al., 2023b; Teknium, 2023), GRAPE curates model-dependent SFT datasets that better matches the base model’s distribution, better mitigating the risks associated with distribution shift like spurious correlations and catastrophic forgetting, while posing minimum overhead.
+This is achieved by calculating the probability of each response using the target model and selects the one with the highest length-normalized probability(§3).After obtaining these more “in-distribution” responses, GRAPE proceeds with standard SFT without any modification to the training.
+Unlike existing datasets that usually contains one-size-fits-all responses for each instruction without customization (Yu et al., 2024; Yuan et al., 2024b; Lian et al., 2023b; Teknium, 2023), GRAPE curates model-dependent SFT datasets that better matches the base model’s distribution, better mitigating the risks associated with distribution shift like spurious correlations (Zhou et al., 2024d) and catastrophic forgetting (Luo et al., 2025; Kotha et al., 2024), while posing minimum overhead of single forward pass over the candidate set.
+In return, GRAPE allows better downstream performance with reduced training compute.
 To GRAPE’s advantage, many existing datasets share overlapping instructions but contain different high-quality responses, e.g., the instructions in Flan (Longpre et al., 2023), GSM-8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b), and the post-training recipes that re-use SFT instructions for preference learning (Lambert et al., 2024; OLMo et al., 2025).
-Therefore, GRAPE can directly select, for each model, the best fit(s) among the off-the-shelf responses without producing new responses, which we do in the experiments(§ 5).
-We first validate our hypothesis through extensive controlled experiments on a reasoning dataset - UltraInteract-SFT (Yuan et al., 2024) and demonstrate the importance of supervising base models with in-distribution responses(§4).
-We experimented on 4 popular pretrained LMs from Mistral (MistralAI, 2024c), Llama3.1 (Dubey et al., 2024) and Qwen2.5 (Hui et al., 2024) families.
-Notably, models fine-tuned with GRAPE-selected responses outperform those trained on a 3$\times$ larger datasets up to $17.3\%$ absolute gain on average performances, and responses from the strongest model under consideration - Llama3.1-405B-Instruct by significant levels.
-We then experiment in realistic scenarios, collecting responses from post-training data for Tulu3 and Olmo-v2.
-We demonstrate the effectiveness of GRAPE by outperforming training on all these available data that is 4.5 times the size of data curated by GRAPE by up to 6.1%.
-Remarkably, GRAPE allows finetuning a Llama3.1-8B base model to exceed the performance of Tulu-8B-SFT using 1/3 data and half number of epochs.
-Our results prove GRAPE as a simple but effective and scalable approach to improving supervised fine-tuning by aligning training distribution and that of the pre-trained checkpoint.
+Therefore, GRAPE can directly select, for each model, the best fit(s) among the off-the-shelf responses without having to produce new responses, which proves effective in our experiments (§ 5).
+[…]
+We then experiment with a more realistic setting for general-domain instruction-tuning, collecting responses from post-training data for Tulu3 and Olmo-v2.
+Again, GRAPE demonstrates its effectiveness by outperforming state-of-the-art data selection approaches by avg.
+4.6% and a strong baseline trained on all these available data that is 4.5 times larger by up to 6.1%.
+Remarkably, GRAPE allows finetuning a Llama3.1-8B base model to exceed the performance of Tulu-8B-SFT using 1/3 data and half of the epochs.
+Our results reveal that distributional alignment is a crucial and previously underappreciated dimension of effective instruction tuning.
+GRAPE introduces this perspective with a simple, scalable algorithm that consistently outperforms strong baselines with far less data and compute for the actual training.
 2 Background and Motivation
+2.1 Data Engineering for Instruction Tuning
+Data is central to the success of effective instruction tuning,  (Xu et al., 2023; Xia et al., 2024; Chan et al., 2024), featuring both automated data synthesis (Xu et al., 2024a; Zeng et al., 2024; Yu et al., 2024; Wei et al., 2023) and selection (Xia et al., 2024; Chen et al., 2023a; Parkar et al., 2024; Li et al., 2024e).
+Some selection approaches focus on high-quality data by leveraging LLMs (Chen et al., 2023a; Parkar et al., 2024; Li et al., 2024b) or employing principled metrics (Kang et al., 2024; Mekala et al., 2024; Xia et al., 2024), while others, such as  Yang et al.
+(2024b); Das & Khetan (2023), aim to identify diversity-optimized subsets for greater efficiency.
+An emerging trend is the customization of training data based on the characteristics of the base models.
+(Li et al., 2024c; Du et al., 2023; Li et al., 2024a).
+These methods typically reweight or filter instructions.
+However, a key overlooked aspect is response selection—specifically, choosing responses that align with the model’s pretrained distribution, which may be critical for preserving useful behaviors and ensuring effective fine-tuning.
+Figure 1: Average performance curve of Llama3.1-8B on Tulu-3-Olmo-2 combo of GRAPE against random selection
+2.2 Toward Distribution-Aligned Supervised Fine-Tuning
 An Analogy from Reinforcement Learning and Preference Learning
 The investigation of this work into the distribution match between the pre-trained LM and supervised fine-tuning (SFT) data is inspired by recent findings on policy optimization for LM alignment with RL (Ouyang et al., 2022) and preference learning (Rafailov et al., 2023; Ethayarajh et al., 2024).
-While the importance of matching training data distribution with the policy has been well noted in both traditional RL (Shi et al., 2023; Fujimoto et al., 2018; Kumar et al., 2019; Peng et al., 2019; Wang et al., 2021; Arora & Goyal, 2023; Jiang & Li, 2016; Tang & Abbeel, 2010) and LM settings (Xiong et al., 2024), preference learning algorithms like DPO (Rafailov et al., 2023), IPO (Azar et al., 2023b) and KTO (Ethayarajh et al., 2024) first emerged as off-policy algorithms.
-However, subsequent research has highlighted the performance gap between on-policy and off-policy training due to distribution shifts  (Xu et al., 2024b; Tang et al., 2024b) and proposed various mitigation strategies  (Zhuang et al., 2023; Zhou et al., 2024c; Zhang et al., 2024a; Xiong et al., 2024; Guo et al., 2024b), showing that training models on data more closely aligned with their policy distribution can significantly improve performance, while failing to do so can yield sub-optimal policies or those that are harder to generalize.
+While the importance of matching training data distribution with the policy has been well noted in both traditional RL (Shi et al., 2023; Fujimoto et al., 2018; Kumar et al., 2019; Peng et al., 2019; Wang et al., 2021; Arora & Goyal, 2023; Jiang & Li, 2016; Tang & Abbeel, 2010) and LM settings (Xiong et al., 2024), preference learning algorithms like DPO (Rafailov et al., 2023), IPO (Azar et al., 2023) and KTO (Ethayarajh et al., 2024) first emerged as off-policy algorithms.
+However, subsequent research has highlighted the performance gap between on-policy and off-policy training due to distribution shifts  (xu2024dposuperiorppo; Tang et al., 2024b) and proposed various mitigation strategies  (Zhuang et al., 2023; Zhou et al., 2024c; Zhang et al., 2024a; Xiong et al., 2024; Guo et al., 2024b), showing that training models on data more closely aligned with their policy distribution can significantly improve performance, while failing to do so can yield sub-optimal policies or those that are harder to generalize.
+Works like SPIN (Yuan et al., 2024a) echo the intuition by gradually improving policy through self-play to mitigate distribution shift.
+Figure 2: Training loss curve of Llama3.1-8B on Tulu-3-Olmo-2.
+Best (highest-probability, selected by GRAPE), Random, and Worst (lowest-probability) responses show a clear loss ordering: Best << Random << Worst.
 Hypothesis: Supervised Fine-tuning Benefits From Data That Better Matches Base Distribution
-The base distribution of pre-trained language models—shaped by extensive training on vast and diverse datasets—is inherently robust and generalizable.
+The base distribution of pre-trained language models—shaped by extensive training on vast and diverse datasets—is inherently robust and generalizable (Brown et al., 2020; Saunshi et al., 2021).
 Therefore, during supervised fine-tuning phase, the pre-trained distribution should be carefully preserved (Kumar et al., 2022; Cohen-Wang et al., 2024; He et al., 2023; Yang et al., 2024d; Ding et al., 2023), to best retain the knowledge and capabilities that emerge during pre-training (Zhou et al., 2023).
@@
 We build on principles of on-policy alignment techniques with key distinctions tailored for SFT.
-Given that SFT reprsents a more primitive stage of post-training than preference learning, strictly sampling responses from the base model itself can introduce risks such as instability, bias reinforcement, knowledge stagnation, and overfitting (Herel & Mikolov, 2024; Mobahi et al., 2020; Allen-Zhu & Li, 2023).
-To address this, we advocate for an approach that stays in-distribution while delivering effective supervision to the base model.
+Yet, given that SFT represents an earlier stage of post-training than RL, sampling responses from the base model itself alone can lead to model collapse, as noted in (Shumailov et al., 2024).
+Prior studies have documented similar risks like instability, bias reinforcement, knowledge stagnation, and overfitting (Herel & Mikolov, 2024; Mobahi et al., 2020; Allen-Zhu & Li, 2023; Ghosh et al., 2024; Dong et al., 2025; Zhang et al., 2024d).
+To address this, we advocate for an approach that stays more in-distribution while delivering effective supervision to the base model.
 To this end, we propose to gather and select responses from various sources, and select one that is closest to the target model’s pretrained distribution, which we name GRAPE.
@@
 We introduce GRAPE, a surprisingly simple yet effective methodology to enhance supervised fine-tuning (SFT) by customizing the training data for the base model.
-The key idea is to find responses among a candidate pool for each instruction $x_{i}$ such that they align closely with the base model’s pretrained distribution $\pi_{\theta_{0}}$.
+The key idea is to find a response, among a candidate pool, for each instruction $x_{i}$ that aligns closely with the base model’s pretrained distribution $\pi_{\theta_{0}}$.
 As diagrammed in Figure 3, GRAPE consists of two main steps, followed by standard SFT:
-  1. 1.
-    Response Collection (§3.1) Collect a pool of high-quality candidate responses either from existing datasets or sampling from multiple LLMs.
-  2. 2.
+Response Collection (§3.1) Collect a pool of high-quality candidate responses from various sources.
 Customization (§3.2): For the target model to be finetuned $\pi_{\theta_{0}}$, find the response(s), for each instruction, that are closest to the pretrained distribution of $\pi_{\theta_{0}}$.
 3.1 Collecting Responses from Existing Resources
-For instruction-tuning of language models, high-quality instructions are more difficult to collect than responses (Xu et al., 2024c; Liu et al., 2024a).
+For instruction-tuning of language models, high-quality instructions are more difficult to collect than responses (Xu et al., 2024b; Liu et al., 2024a).
 Therefore, it is a common practice to reuse existing instruction-tuning prompts while generating diverse responses using various methods tailored to specific requirements.
-For instance, instructions from Flan (Longpre et al., 2023), OpenOrca (Lian et al., 2023a), ShareGPT (Team, 2023), and the training splits of GSM-8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b), and CodeContests (Li et al., 2022) are frequently reused in datasets like Olmo (OLMo et al., 2025), Tulu (Lambert et al., 2024), OpenHermes (Teknium, 2023), OpenOrca (Lian et al., 2023a), MetaMath (Yu et al., 2024), MathInstruct (Yue et al., 2023), UltraFeedback (Cui et al., 2024), and UltraInteract (Yuan et al., 2024), whether for SFT or preference learning.
+For instance, instructions from Flan (Longpre et al., 2023), OpenOrca (Lian et al., 2023a), ShareGPT (Team, 2023), and the training splits of GSM-8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b), and CodeContests (Li et al., 2022) are frequently reused in datasets like Olmo (OLMo et al., 2025), Tulu (Lambert et al., 2024), OpenHermes (Teknium, 2023), OpenOrca (Lian et al., 2023a), MetaMath (Yu et al., 2024), MathInstruct (Yue et al., 2023), UltraFeedback (Cui et al., 2024), and UltraInteract (Yuan et al., 2024b), whether for SFT or preference learning.
 The solutions are generated using different models or follow varying styles depending on the specific needs.
@@
 When such resources are unavailable or insufficient, practitioners can generate new responses and apply GRAPE.
-Responses for the $i$-th instruction are collected from various datasets to form $\mathcal{R}_{i}=\{y{{}_{i}^{j}}:j=1,\dots,J^{(i)}\}$, where $J^{(i)}$ denotes the number of responses collected for the $i$-th instruction.
+For each instruction, we collect multiple responses from various datasets. These responses form a candidate set associated with the instruction.
 3.2 Customize Dataset For Models
-We then compute the conditional probability of each response $\pi_{\theta_{0}}(y{{}_{i}^{j}}\mid x_{i})$.
-We rank the responses based on the conditional probability and take those with the highest probability for each instruction.
-Notably, GRAPE’s selection process requires only a forward pass through the candidates and does not require gradient computation.
-Its overhead is substantially lower than many model-based data selection algorithms (Xia et al., 2024; Yang et al., 2024b; Liu et al., 2024b; Zhao et al., 2021; Zhang et al., 2024b; Pan et al., 2024).
-Additionally, it is important to distinguish GRAPE from the other preplexity-based data selection and curriculum planning methods (Wu et al., 2024; Li et al., 2024b; Liu et al., 2024c).
+[…]
+Since GRAPE only involves forward-pass log-probability computation (no gradients or optimization), it is highly efficient and simple to integrate into any SFT pipeline with minimal overhead compared with model-based data selection approaches (Xia et al., 2024; Yang et al., 2024b; Liu et al., 2024b; Zhao et al., 2021; Zhang et al., 2024b; Pan et al., 2024).
+Additionally, it is important to distinguish GRAPE from the other perplexity-based data selection and curriculum planning methods (Wu et al., 2024; Li et al., 2024b; Liu et al., 2024c).
 Existing approaches focus on selecting instructions by using perplexity as a difficulty measure, which differs from GRAPE that uses probability to select for each instruction in a fixed instruction set, responses that better matches with the base model’s distribution.
 Our experiments in §5 demonstrate that low-probability responses with fixed set of instructions are detrimental to performance, further emphasizing the fundamental difference in the two processes.
-4 Preliminary Experiments
-In this section, we seek to empirically verify the hypothesis in §2.
-To evaluate it under a controlled set-up, we consider using instructions from UltraInteract-SFT, and constrain within chain-of-thought reasoning (predominantly for coding, logic and math.) scenario.
-This allows us to control confounding factors like style and format while ensuring manageable data quality control during data collection.
-The insights from this experiment will inform the later more realistic setting in §5.
+4 Controlled Experiments Show the Benefits of Distributional Alignment
+We conduct controlled experiments using UltraInteract-SFT to test whether selecting responses aligned with a base model’s distribution improves fine-tuning outcomes more than relying on stronger generators or scaling data size.
+This setup—focused on verifiable tasks in coding, logic, and math—isolates the effect of distribution matching.
+Results show that GRAPE-selected responses consistently outperform alternatives, validating distributional alignment as a key supervision signal.
+These findings motivate our broader evaluations in §5.
 4.1 Experimental Setup
-We detail our experiment settings below.
-4.1.1 Training Data Curation
-In this controlled experiment, we focus on chain-of-thought reasoning (Wei et al., 2022; Wang et al., 2024; Luo et al., 2024; Cobbe et al., 2021; Li et al., 2023; Lightman et al., 2023).
+Training Data Curation
+In this controlled experiment, we focus on chain-of-thought reasoning (Wei et al., 2022; Wang et al., 2024a; Luo et al., 2024; Cobbe et al., 2021; Li et al., 2023; Lightman et al., 2023).
 Different models may follow different reasoning paths to solve a problem, while their final solutions can be easily verified.
-Figure 2: The training loss curve of Llama3.1-8B on Tulu-3-Olmo-2 combo.
-Best stands for training with highest-probability responses (GRAPE-selected); Random stands for randomly chosen responses for each instruction in the pool, Worst stands for lowest probability responses in the base model’s distribution.
-Throughout training, Best has lower loss than Random than Worst.
-We use UltraInteract-SFT (Yuan et al., 2024), which contains approximately $80,800$ unique instructions covering coding, math (chain-of-thought and program-aided) and logic reasoning domains , where each instruction is paired with varying numbers of different responses to contain a total $>280,000$ training examples.
+We use UltraInteract-SFT (Yuan et al., 2024b), which contains approximately $80,800$ unique instructions covering coding, math (chain-of-thought and program-aided) and logic reasoning domains , where each instruction is paired with varying numbers (avg.
+3.5/instruction) of different responses to contain a total >280,000>280,000 training examples.
 The responses in the dataset are strictly in step-wise format.
-For each instruction, we use GRAPE to select the top-ranking responses, ensuring that the number of responses matches the original UltraInteract-SFT dataset for fair comparisons.
-Response Generation
+For each instruction, we construct a response pool consisting of both (i) original UltraInteract-SFT responses and (ii) additional responses generated from a diverse set of LLMs.
+GRAPE is then applied to select the most in-distribution response per instruction to this enlarged candidate pool and ensure the number of responses matches the original UltraInteract-SFT dataset for fair comparisons.
 We collect responses from a diverse set of models of various sizes across model families, including Mixtral-7x7B-Instruct (Jiang et al., 2024), Codestral-22B (MistralAI, 2024a), Mistral-Small (MistralAI, 2024b), Llama-3.1-70B-Instruct and Llama-3.1-405B-Instruct (Dubey et al., 2024), and Qwen2.5-72B-Chat (Yang et al., 2024a), resulting in approximately 10x additional responses per instruction.
 The responses are then filtered based on the answers to ensure their validity following Yuan et al.
-(2024).
-Figure 3: The average performance curve of Llama3.1-8B on Tulu-3-Olmo-2 combo.
-Best stands for training with highest-likelihood responses (GRAPE-selected); Random stands for randomly chosen responses for each instruction in the pool.
-The average is taken over all items on table 2 except AlpacaEval2.
-4.1.2 Base Models
+(2024b).
+Base Models
 To demonstrate the generalizability of GRAPE, we evaluate its performance across multiple LLMs, including Llama-3.1-8B and Llama-3.2-3B from Llama-3 (Grattafiori et al., 2024) family, Mistral-7B (Jiang et al., 2023) and Qwen2.5-7B (Hui et al., 2024).
-4.1.3 Evaluation
+We ensure that all training configurations (GRAPE, baselines) use the same number of instructions and responses per instruction as original UltraInteract-SFT, unless otherwise stated.
+Evaluation
 We evaluate the model on coding and math reasoning benchmarks.
@@
 MATH contains high-school level math competition problems, whereas GSM-Plus is a more challenging variant of GSM-8k (Cobbe et al., 2021) and Theorem-QA contains complex math reasoning problems.
-4.1.4 Baselines
-We compare GRAPE with several baselines.
-Except for Up-scaled Dataset which is 3×3$\times$3 × larger, the rest are controlled for the number of responses per-instruction to be the same as UltraInteract-SFT.
-  • Original dataset. Performing SFT over the original UltraInteract-SFT dataset as is. The original dataset also contains high-quality, verified responses.
-    We compare GRAPE to directly using a standard, well-curated dataset.
-Responses from a powerful model are likely to represent an upper-bound quality of training data, and comparing against this baseline allows us to isolate whether further improvements stem from the strategic response selection performed by GRAPE.
-  • Responses from the strongest model under consideration. We use responses exclusively from the best model used to generate responses (in our case, Llama3.1-405B-Instruct).
-  • Up-scaled Dataset.
-The comparison with a larger dataset assesses how well selection improves instruction tuning compared to training on a larger set of diverse correct responses.
-Each instruction is paired with three times of randomly selected responses from the pool.
-[^1] For example, if UltraInteract contains 3 different responses for instruction $x$, we end up including 9 for this setting.
-We train all models for 1 epoch with a learning rate of $10^{-5}$.
+Baselines
+[…]
 MATH GSMPlus TheoremQA
@@
 3x Data 48.1 13.3 52.8 25.1 26.1 50.0 49.4 16.8 12.4 32.7 3.6
-Mistral-7B Ours 52.4 15.6 53.4 28.9 34.6 50.5 52.8 17.8 20.6 36.3 -
+Mistral-7B GRAPE 52.4 15.6 53.4 28.9 34.6 50.5 52.8 17.8 20.6 36.3 -
 Original-UI 54.3 11.1 58.9 29.7 31.0 53.7 51.6 20.0 20.8 36.8 4.7
@@
 3x Data 48.8 7.8 57.9 25.5 11.2 48.6 45.1 20.6 19.6 31.7 9.8
-Llama3.1-8B Ours 57.3 19.4 63.8 34.8 39.2 56.6 56.1 22.5 23.9 41.5 -
+Llama3.1-8B GRAPE 57.3 19.4 63.8 34.8 39.2 56.6 56.1 22.5 23.9 41.5 -
 Original-UI 32.9 3.9 41.6 12.8 16.1 30.8 19.5 14.6 10.5 20.3 3.8
@@
 3x Data 42.6 6.7 42.9 8.7 5.1 17.8 19.5 14.6 12.8 19.0 5.1
-Llama3.2-3B Ours 42.6 13.3 44.6 16.4 17.6 34.9 20.6 15.1 11.4 24.1 -
+Llama3.2-3B GRAPE 42.6 13.3 44.6 16.4 17.6 34.9 20.6 15.1 11.4 24.1 -
 Original-UI 67.0 41.2 60.0 51.0 38.3 64.1 59.5 14.6 10.5 45.1 11.6
@@
 3x Data 75.6 48.3 62.9 32.8 24.5 47.1 22.8 21.6 19.0 39.4 17.3
-Qwen2.5-7B Ours 77.4 48.9 70.7 56.4 45.3 67.7 66.3 37.4 40.1 56.7 -
-Table 1: Result of synthetic experiment on UltraInteract-SFT. The last column, Abs. $\Delta$ is the absolute improvement of GRAPE over that row.
+Qwen2.5-7B GRAPE 77.4 48.9 70.7 56.4 45.3 67.7 66.3 37.4 40.1 56.7 -
+Table 1: Result of synthetic experiment on UltraInteract-SFT. The last column, Abs. $\Delta$ is GRAPE’s absolute improvement on average performance over that row.
 4.2 Results and Analysis
@@
 The details are presented below.
+We discuss additional results in Appendix B
 5.1 Data Mixture Details
@@
 To demonstrate the effectiveness of GRAPE, we collected the overlapping instructions from both models and gather their corresponding responses.
+Figure 4: Breakdown of GRAPE-selected responses for 1K Tulu instructions vary significantly across base models, reflecting its highly model-oriented nature over responses. Details in Appendix H
 From the preference data, we retained only the winning responses.
@@
 We evaluated on LeetCode (Guo et al., 2024a), MATH (Hendrycks et al., 2021b), BigBenchHard(BBH) (Suzgun et al., 2022), MMLU (Hendrycks et al., 2021a), and AlpacaEval-V2 (Dubois et al., 2024).
-LeetCode, MATH, BBH and MMLU are evaluated the in the same way as in (Yuan et al., 2024), where we use zero-shot for MATH and MMLU, 3-shot example for BBH.
+LeetCode, MATH, BBH and MMLU are evaluated the in the same way as in (Yuan et al., 2024b), where we use zero-shot for MATH and MMLU, 3-shot example for BBH.
 We use the same AlpacaEval-v2 as in OpenInstruct.
 5.3 Baselines
-  • Responses From The Original SFT on The Subset. In this experiment, we keep the same set of instructions (350.4k) and pair each instruction with the response from the SFT mixture.
-  • Random Response Candidate From The Pool.
-Randomly pairing each instruction with a candidate response for that instruction, which may come from either the SFT mixture or a winning response in the preference-learning data.
-  • Lowest-Probability Response. Instead of taking the highest probability response to each instruction for the base model, we do the opposite on and take the lowest for each.
-  • Entire SFT Dataset. We train on the entire SFT data mixture of Tulu with 939k instances.
-  • All Available Responses For The Overlapping Instructions. In this setting, we train on the entire candidate pool for all overlapping instructions, resulting in 103.6k instances.
-  • S2L (Yang et al., 2024b).
-S2L is a state-of-the-art unsupervised data selection baseline designed to select balanced subsets from large datasets by leveraging training loss trajectories.
-It first trains a small reference model within the same model family and record loss trajectory for each training example, then applies K-means clustering on the loss trajectory and samples equally from each.
-We use Llama-3.2-1B as reference for Llama3.1-8B and Qwen2.5-0.5B for Qwen2.5-7B.
-For Mistral-v0.3-7B, we use itself as reference since no smaller models in the family are available.
-For S2L, we select the same number of data from the same pool as GRAPE.
-Further details in Appendix B.
-  • All data available for SFT mixture.
-We train with all the available responses from all these datasets that share prompts with the SFT mixture.
-Which gives approximately 1.58 million (4.5 times thesize of GRAPE-curated dataset).
-This is the union of the entire dataset and GRAPE’s candidate pool.
+We extensively compare against three baseline types: controlled baselines with fixed instructions, scaling baselines with increased data, and state-of-the-art selection methods.
+Additional details in Appendix C
+Controlled Baselines
+We include three baselines to isolate the effect of GRAPE’s response selection.
+SFT-only replaces GRAPE-selected responses with those from the original SFT dataset, pairing each instruction with a standard reference response to measure improvement over presumably good SFT responses.
+Random selects candidate responses uniformly at random from the pool for the same set of instructions, establishing a noise-tolerant baseline.
+Reverse-GRAPE instead selects responses with the highest perplexity.
+We test if responses diverging the most from the base model’s distribution degrade performance, providing a contrast that sharpens the effectiveness of GRAPE.
+Scaling Baselines
+[…]
+State-Of-The-Art SFT Data Selection Approaches
+We compare GRAPE against recent state-of-the-art data selection methods.
+LESS (Xia et al., 2024) selects data based on influence scores on validation tasks.
+We follow the implementation setup from Dai et al.
+(Dai et al., 2025).
+Emb-NV selects training data close to validation in embedding space using NV-Embed-V2 (Lee et al., 2025), following (Ivison et al., 2025).
+S2L (Yang et al., 2024b) clusters data via loss trajectories from small reference models (Llama-3.2-1B, Qwen2.5-0.5B, Mistral-v0.3-7B) and samples uniformly across clusters to match GRAPE’s data budget.
 AlpacaEval2
@@
 S2L 350.4k 8.5 7.6 68.6 63.1 26.5 16.1 31.7 4.2
+Emb-NV 350.4k 8.4 7.0 67.9 63.7 32.1 17.2 32.7 3.2
+LESS 350.4k 7.3 6.0 68.2 63.3 25.1 16.1 31.0 4.0
 All Available Data 1.58M 8.8 10.1 69.8 62.1 32.5 16.1 33.2 2.7
@@
 S2L 350.4k 10.5 11.9 61.9 57.1 22.4 13.9 29.6 2.3
+Emb-NV 350.4k 6.0 5.2 60.7 56.5 23.9 11.7 27.3 4.6
+LESS 350.4k 6.4 4.8 59.2 55.4 16.0 8.3 25.0 6.9
 All Available Data 1.58M 8.0 7.0 55.3 53.7 25.4 12.3 26.9 5.0
@@
 S2L 350.4k 13.4 14.9 72.7 73.1 53.4 40.6 44.7 3.9
+Emb-NV 350.4k 11.9 10.6 72.1 72.5 53.3 43.3 44.0 4.6
+LESS 350.4k 7.8 5.9 71.3 72.9 48.3 41.1 41.2 7.4
 All Available Data 1.58M 13.3 12.3 70.3 71.8 44.0 42.8 42.4 6.1
 Qwen2.5 -7B GRAPE 350.4k 20.0 20.4 73.2 73.3 60.0 44.4 48.6 -
-Table 2: GRAPE on the Tulu-Olmo collection.
-For Llama3.1-8B base model, we included Tulu3-SFT model’s results.
-The “*”-marked number for MATH is obtained using 4-shot prompting.
-We train all the models (except that we took Tulu3-SFT numbers directly) for 1 epoch with a learning rate of $10^{-5}$.
+[…]
 5.4 Results
+GRAPE
+Qwen2.5
+-72B
+Llama3.1
+-405B
+Gemma
+-it-9B
+LC 28.1 25.8 16.3 26.9
+WR 33.1 24.3 17.0 20.2
+Table 3: Alpaca-Eval2 On Magpie-Zoo Xu et al. (2024c).
 As shown in Table 2, models fine-tuned on responses selected by GRAPE outperforms the strong baselines we constructed, especially the one that trains over all available data by significant margins across the 3 models.
 Remarkably, using roughly 1/6 training computation (Tulu3-8B-SFT was trained for 2 epochs on 3 times of data), our performance exceeds that of Tulu3-8B-SFT.
-Also, GRAPE outperforms state-of-the-art data-selection approaches like S2L, despite its simplicity and efficiency, further highlighting its effectiveness in diverse real-world scenarios and strengthening its overall practicability for large-scale instruction tuning and robust generalization.
+Also, GRAPE outperforms state-of-the-art data-selection approaches like S2L, despite its simplicity and efficiency, further highlighting its effectiveness in diverse real-world scenarios and making it a practical option for real-world SFT setups with minimal data engineering effort.
 Without the need to synthesize any new data, one can easily leverage established datasets sourced from the web to customize a dataset for each base model that yields better fine-tuning outcome.
-These results feature GRAPE not only as an effective strategy to enhance performances, but a handy approach to improve fine-tuning efficiency.
-5.5 Further Experiments
-Item
-Metric Data
+These results highlight GRAPE as an effective and efficient selection strategy for real-world SFT.
+5.5 Ablations
+Remains Effective Even with a Single Response Source
+In Section 4, we demonstrated how GRAPE enables practitioners to refine model-generated responses for improved training outcomes.
+Beyond that, GRAPE can optimize responses from a single generator.
+The Magpie-Zoo (Xu et al., 2024c) dataset contains a fixed set of instructions and multiple versions of response sets each generated by different language models.
+Using the Magpie-Zoo instruction set, we sample 10 responses per instruction from Qwen2.5-72B-Instruct, select in-distribution responses with GRAPE, to train a Mistral-v0.3-7B model.
+We compare with replicas that achieve top-3 performance from Magpie-Zoo.
+Model Data Avg.
+Self-Distilled 28.4(-)
+Mistral-7B Original-UI 32.7
+Self-Distilled 29.4(-)
+Llama3.1-8B Original-UI 36.8
+Self-Distilled 15.1(-)
+Llama3.2-3B Original-UI 20.3
+Table 4: Performance Degradation From Self-Distillation On UI.
+As shown in Table 3, GRAPE-selected responses further boost the performance.
+Given that batch sampling from a strong model introduces minimal latency (Zhong et al., 2024; Zhou et al., 2024e), this result positions GRAPE as a practical and efficient data curation strategy—achieving strong results with just a single generator and no additional engineering overhead.
+Why GRAPE Outperforms Self-Generated Responses
+Dataset Model Response Generator N Acc.
+MATH Mistral Llama3.1-70B 10 18.2
+MATH Mistral FT-Mistral 10 15.9 (-)
+MATH Llemma Llama3.1-70B 10 26.2
+MATH Llemma FT-Llemma 10 23.6 (-)
+MATH Mistral $M$-AnsAug 22.3
+MATH Mistral FT-Mistral 10 20.6 (-)
+MATH Llemma $M$-AnsAug 28.1
+MATH Llemma FT-Llemma 10 21.4 (-)
+Table 5: Self-generation on MATH dataset.
+FT-Mistral refers to the model right above that row finetuned from either $M$-AnsAug or Llama3.1-70B-Instruct produced solutions.
+N stands for the number of responses sampled.
+To probe the effectiveness of GRAPE, we ablate it against a degenerate alternative: self-generation, where the model is fine-tuned on its own outputs.
+We fine-tune a base model using responses generated by its previously fine-tuned variant.
+This setup consistently degrades performance (Table 4).
+We further confirm this on the MATH dataset (Hendrycks et al., 2021b): responses from strong models like Llama3.1-70B-Instruct or MetaMathQA (Yu et al., 2024) are used to fine-tune a model that then generates new solutions for another round of fine-tuning.
+Again, performance drops (Table 5).
+This failure arises from distributional collapse (see § 2.2): self-generated responses become increasingly narrow and repetitive, reinforcing biases and reducing exposure to diverse reasoning.
+Correctness alone is insufficient—external diversity is essential for generalization.
+GRAPE avoids this collapse by selecting external responses that are both diverse and distribution-aligned, preserving semantic breadth and stylistic variability while staying true to the model’s pretraining.
+This enables more stable and generalizable fine-tuning.
+6 Conclusion
+We present GRAPE, a simple yet effective method for improving supervised fine-tuning by selecting responses aligned with the base model’s pretrained distribution.
+GRAPE requires only a forward pass over candidate responses, making it highly efficient and easy to integrate.
+Despite its simplicity, GRAPE consistently outperforms stronger baselines using significantly larger datasets and surpasses more complex, costly data selection methods.
+Our study affirms that carefully aligning SFT data with a model’s pretrained distribution yields substantial performance and efficiency gains.
+7 Discussion and Limitations
+Response versus Instance Level Selection
+GRAPE selects responses: it begins with a fixed set of instructions and evaluates multiple candidate responses for each.
+This differs from instance-level selection approaches (e.g., Kung et al.
+(2023); Li et al.
+(2024c); Wang et al.
+(2024b)), which focus on choosing which instructions to include based on factors such as coverage, skill-balancing or difficulty.
+In contrast, GRAPE focuses on the quality of supervision—that is, selecting responses to provide the most effective learning signal.
+Importantly, this response-centric perspective is complementary, not contradictory, to instance-level or instruction-based selection methods; both might be combined to enhance overall data quality and training efficiency.
+Limitations
+Like many other data selection algorithms (Du et al., 2023; Li et al., 2024c; Xia et al., 2024; Das & Khetan, 2023; Kang et al., 2024; Mekala et al., 2024; Yang et al., 2024c; Zhang et al., 2024b; Pan et al., 2024; Dai et al., 2025; Ivison et al., 2025; Bhatt et al., 2024; Yin & Rush, 2024; Liu et al., 2024b), GRAPE assumes a quality-controlled candidate pool from which to select samples.
+Furthermore, because GRAPE relies on the base model itself, its selection effectiveness may be influenced by the model’s inherent capabilities.
+8 Acknowledgment
+This project is partly supported by NSF under award No.
+2019897.
+This research used the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois.
+DeltaAI is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: See experiments in Section 4 and 5.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: See section 5.6, where we show pursuing in-distribution answers in the wrong way can lead to performance degradations.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate ”Limitations” section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: We do not have theoretical results.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: See appendices.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24.
+Justification: We use publicly available datasets and models.
+Our method only requires computing the normalized probability of training data, which can be easily done with any open-sourced machine learning codebase.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: See Experiments Sections and Appendices.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: We follow standard evaluation paradigms by computing pass@1 accuracy with greedy sampling, which does not explicitly involve randomness.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer ”Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: See appendicies, where we provide information about GPU resources.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: See paper.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49.
+Justification: This paper presents work whose goal is to advance the field of Machine Learning.
+It investigates fundamental aspects of instruction-tuning of language models and should not have direct societal impacts or implications that should be discussed here specifically, to the best of the authors’ knowledge.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: We do not release data or models that have a high risk for misuse.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: Yes, we cite all the datasets and pretrained models used in the paper, which are all open-sourced for research use.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: We do not release new assets.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: We do not have crowdsourcing experiments or research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: We do not have crowdsourcing experiments or research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: We do not involve LLMs as a core component in implementing or developing the method of this work.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Appendix A Parameter Distance
+We measure the L2 norm of model parameter difference between fine-tuned and pre-trained checkpoints, as a signal of how much the distribution has drifted during SFT ichi Amari (2016); Cover & Thomas (2006).
+We notice that training over well-matched distribution shifts the parameter less than training over those ill-matched.
+Mistral-7B-v0.3 Llama3.1-8B Qwen2.5
+GRAPE 8.006 8.196 8.426
+Worst 8.029 8.202 8.467
+Table 6: Performance comparison across different models
+Appendix B Further Experiments
+B.1 Comparing with Reward Based Selection
+We compare GRAPE with purely reward-based selection.
+where for each instruction, we select the response with the highest scalar reward as determined by a reward model - Skywork-Reward-Llama3.1-8B-v0.2.
+Once the top-ranked response for each instruction is selected, we proceed with standard supervised fine-tuning on the resulting instruction-response pairs.The RFT setup provides a natural contrast to our proposed GRAPE method by emphasizing reward alignment over base-model alignment, thereby enabling us to disentangle the effects of distribution matching versus reward optimization in SFT data selection.
+As shown in Table 7, GRAPE outperforms reward-based selection across both models and all benchmarks.
+These results suggest that aligning supervision with the base model’s own distributional preferences—rather than relying on external reward models—can yield better task performance.
+Model Method AE WR AE WR (LC) LeetCode MATH MMLU BBH Avg
+LLaMA3.1-8B GRAPE 15.2 14.8 19.4 32.1 64.5 69.6 35.9
+Reward 14.0 14.5 17.2 31.3 63.3 69.0 34.9
+Mistral-v0.3-7B GRAPE 13.9 13.6 18.3 24.2 59.2 62.3 31.9
+Reward 12.5 13.9 13.3 22.4 58.5 62.2 30.5
+Table 7: Performance comparison between GRAPE and reward-based selection across benchmarks for LLaMA3.1-8B and Mistral-v0.3-7B. Metrics are benchmark-specific scores (higher is better).
+B.2 Experiment on OpenHermes
+Item Metric Data
 Llama
@@
 Table 8: Results on OpenHermes-2.5. The Subset row refers to training exclusively on the SFT responses over the subset.
-We extend our evaluation using OpenHermes-2.5 (Teknium, 2023), a high-quality, open instruction-tuning dataset containing approximately 1 million distinct instructions.
-Following a similar strategy to the previous combo experiment, we source responses from additional datasets from  Huang et al.
+To test the generality of our findings beyond the UltraInteract and Tulu-Olmo settings, we conduct additional experiments on the OpenHermes-2.5 (Teknium, 2023) dataset—a large-scale, high-quality instruction-tuning corpus with approximately 1 million distinct instructions.
+Following the setup from §5, we apply GRAPE to select from responses aggregated across sources, including Huang et al.
 (2024) and HuggingFace-H4 (2024).
-For preference-based datasets, we include only the winning responses, consistent with the methodology in the Tulu-Olmo experiment.
-This process results in 575K unique instructions and 1.34 million total instruction-response pairs.
-As shown in Table 3, GRAPE-selected data yields better performances too.
-The consistent improvement across models reaffirm our earlier observations: GRAPE serves as an adaptive response selection mechanism that significantly enhances SFT performance that has wide applicability.
-5.6 Discussion:GRAPE works if all responses are from the same LM
-GRAPE
-Qwen2.5
--72B
-Llama3.1
--405B
-Gemma
--it-9B
-LC 28.1 25.8 16.3 26.9
-WR 33.1 24.3 17.0 20.2
-Table 4: Alpaca-Eval2 Results On Magpie-Zoo (Xu et al., 2024d).
-In Section 4, we demonstrated how GRAPE enables practitioners to refine model-generated responses for improved training outcomes.
-Beyond that, GRAPE can optimize responses from a single generator.
-Using the Magpie-Zoo instruction set (Xu et al., 2024d), we sample 10 responses per instruction from Qwen2.5-72B-Instruct, select in-distribution responses with GRAPE, to train a Mistral-v0.3-7B model.
-We also compare with training on top-3 reported datasets of Magpie-Zoo.
-As shown in Table 4, GRAPE-selected responses further boost the performance.
-Practically, batch sampling of responses cause little latency (Zhong et al., 2024; Zhou et al., 2024d).
-One therefore can sample batches of responses from a single strong model and apply GRAPE on top of it to find the response closest to the base model’s distribution efficiently.
-5.7 Discussion: Is In-Distribution A Silver Bullet?
-The success of GRAPE verifies our key hypothesis that matching SFT distribution to the base model benefits the performance.
-Yet, we argue that this cannot be pushed to the limit of SFT using data from the same distribution.
-By experimenting with training on solutions sampled from a trained version of the same base model, we notice a drastic performance drop from the original model as shown in Table 5.
-Model Data Avg.
-Self-Distilled 28.4(-)
-Mistral-7B Original-UI 32.7
-Self-Distilled 29.4(-)
-Llama3.1-8B Original-UI 36.8
-Self-Distilled 15.1(-)
-Llama3.2-3B Original-UI 20.3
-Table 5: Performance Degradation From Self-Distillation On UI.
-We explored this using the MATH (Hendrycks et al., 2021b) dataset.
-A base model was fine-tuned using solutions from stronger models like Llama3.1-70B-Instruct or GPT-3.5-Turbo-augmented MetaMathQA (Yu et al., 2024).
-This fine-tuned model then generated new solutions, which were used to further fine-tune another base model instance, simulating iterative on-policy tuning.
-Dataset Model Response Generator N Acc.
-MATH Mistral Llama3.1-70B 10 18.2
-MATH Mistral FT-Mistral 10 15.9 (-)
-MATH Llemma Llama3.1-70B 10 26.2
-MATH Llemma FT-Llemma 10 23.6 (-)
-MATH Mistral MM-AnsAug - 22.3
-MATH Mistral FT-Mistral 10 20.6 (-)
-MATH Llemma MM-AnsAug - 28.1
-MATH Llemma FT-Llemma 10 21.4 (-)
-Table 6: Effect of self-generation on MATH dataset.
-FT-Mistral refers to the model right above that row finetuned from either MM-AnsAug or Llama3.1-70B-Instruct produced solutions.
-N stands for the number of responses sampled.
-As shown in Table 6, performance declined when solutions were sampled from the fine-tuned model itself.
-This decline stems from reduced solution diversity, leading to distributional drift and poorer generalization.
-These results validate our claim in §2 that robust fine-tuning requires more than alignment with the base model—it needs complementary strategies to maintain response diversity and quality.
-6 Related Works
-Data Engineering For Instruction Tuning
-Data is central to the success of effective instruction tuning,  (Xu et al., 2023; Xia et al., 2024; Chan et al., 2024), featuring both automated data synthesis (Xu et al., 2024a; Zeng et al., 2024; Yu et al., 2024; Wei et al., 2023) and selection (Xia et al., 2024; Chen et al., 2023a; Parkar et al., 2024; Li et al., 2024e).
-Some selection approaches focus on high-quality data by leveraging LLMs (Chen et al., 2023a; Parkar et al., 2024; Li et al., 2024b) or employing principled metrics (Kang et al., 2024; Mekala et al., 2024; Xia et al., 2024), while others, such as  Yang et al.
-(2024c); Das & Khetan (2023), aim to identify diversity-optimized subsets for greater efficiency.
-Another emerging trend is the customization of training data based on the characteristics of the base models.
-For instance,  Li et al.
-(2024c); Du et al.
-(2023) leverage the base model itself to select a subset of instructions, while  Li et al.
-(2024a) introduce a teacher model to guide the selection process.
-However, these approaches primarily focus on identifying a re-weighted subset of questions for training.
-In contrast, GRAPE has a different focus, where it aims to find a set of responses from the solution space that not only provide good coverage of the golden distribution but also align closely with the base model’s policy.
-On-Policy Methods For Language Model Alignment
-Recent advances in RLHF (Ouyang et al., 2022), RLAIF (Lee et al., 2024b) and preference learning (Rafailov et al., 2023) research have identified the benefits of on-policy data for model alignment (Xu et al., 2024b; Tajwar et al., 2024; Zhang et al., 2024c; Guo et al., 2024b; Liu et al., 2024d; Zhang et al., 2024a; Miao et al., 2024; Gulcehre et al., 2023).
-However, these works primarily emphasize the preference-tuning stage, where models are refined to generate samples from an already reasonable policy.
-In contrast, less attention is given to the supervised fine-tuning phase, which starts with base models that are not yet well-prepared to generate high-quality answers that align sufficiently with the desired policy for optimization.
-7 Conclusion
-We present GRAPE, a simple yet highly effective approach to improve supervised fine-tuning data building on our empirically-verified hypothesis that instruction tuning data shall better match the base model’s distribution to optimize the training outcome.
-GRAPE is surprisingly simple and efficient, where it customizes response data to supervise each model by taking highest-probability instances.
-We demonstrated the effectiveness of GRAPE on multiple settings.
-Remarkably, GRAPE can outperform not only SOTA data selection baselines, but also Tulu3-SFT by using a 1/3 subset of its SFT instructions, and performs better than training on 4.5 times data it selects from.
-We show GRAPE as a highly scalable and simple approach can be adopted in multiple practical scenarios with minimal overhead.
-Appendix A Parameter Distance
-We measure the L2 norm of model parameter difference between fine-tuned and pre-trained checkpoints, as a signal of how much the distribution has drifted during SFT (ichi Amari, 2016; Cover & Thomas, 2006).
-We notice that training over well-matched distribution shifts the parameter less than training over those ill-matched.
-Mistral-7B-v0.3 Llama3.1-8B Qwen2.5
-GRAPE 8.006 8.196 8.426
-Worst 8.029 8.202 8.467
-Table 7: Performance comparison across different models
-Appendix B Further Details On S2L
-This section details the experimental setup for two unsupervised data selection baseline, S2L applied to the joined pool of Tulu-3 and Olmo-v2 data.
+For preference-based datasets, we retain only the winning responses to ensure quality, mirroring our earlier selection protocol.
+This results in 575K unique instructions and 1.34M instruction-response pairs.
+As shown in Table 8, GRAPE continues to outperform naive combination strategies.
+The consistent gains across diverse data sources and model families strengthen our central claim: GRAPE is a general-purpose, model-aligned response selection strategy that reliably improves SFT performance in real-world, large-scale instruction tuning.
+B.3 Data Selection For Long Chain-of-Thoughts
+O1-/R1-style long chain-of-thoughts have drawn increasing attention.
+This paradigm, exemplified by models like OpenAI’s O1 and DeepSeek-R1, has shown remarkable success in challenging domains such as mathematics and coding.
+We further experiment with the use of GRAPE in long chain-of-thought distillation.
+We generate multiple candidate trajectories using R1-Distill-Qwen-1.5B DeepSeek-AI et al.
+(2025) for a subset of OpenR1-Math (Face, 2025) dataset — LUFFY (Yan et al., 2025), and verify the correctness of each, retaining the correct ones.
+We compare the results on lowest versus highest perplexity instances below in table 9.
+Model Perplexity Acc.
+Qwen2.5-1.5B Highest 0.330
+Qwen2.5-1.5B Lowest (GRAPE) 0.396
+Qwen2.5-3B Highest 0.524
+Qwen2.5-3B Lowest (GRAPE) 0.544
+Table 9: Performance metrics on MATH dataset
+B.4 Token-level GRAPE
+To further investigate the alternative uses of our insight, we conduct experiments beyond data selection by incorporating token-level likelihoods directly into the training objective.
+Specifically, we modified the loss function to weigh each token proportionally to its likelihood.
+We trained this variant on the OpenThoughts‑114k Team (2025) dataset, a curated collection of 114k high-quality reasoning samples spanning domains such as math, science, code etc.
+Evaluation followed the same benchmark suite as LUFFY Yan et al.
+(2025), including competition-level math datasets (AIME24/25, AMC, MATH‑500, Minerva, OlympiadBench) and general reasoning tests (ARC‑c, GPQA‑Diamond, MMLU‑Pro).
+Our result in Table 10 show that our token-level likelihood-weighted training yields consistent improvements across several benchmarks.
+Benchmark Baseline (%) Token-level GRAPE (%)
+MATH 55.60 60.20
+OlympiadBench 21.04 28.30
+Minerva 15.07 19.85
+AIME-24 3.12 4.90
+AMC 23.76 29.74
+AIME-25 4.38 3.23
+ARC-C 60.84 76.62
+GPQA-Diamond 7.07 19.70
+MMLU-Pro 27.02 34.35
+Table 10: Qwen2.5-3B performance (in %) on various benchmarks under Baseline vs. Token-level GRAPE.
+Appendix C Further Details On Baselines
+This section details the experimental setup for our data selection baselines: S2L,LESS and NV-Embed.
+C.1 S2L
 S2L, a state-of-the-art unsupervised data selection baseline, operates through two key steps: training a reference model to capture training dynamics and clustering the resulting trajectories to form a diverse, balanced subset of training data.
@@
 From each cluster, an equal number of examples are sampled to maintain a balanced subset distribution.
-Appendix C Further Training Details
+C.2 LESS
+LESS is a state-of-the-art model-based and supervised data selection method that leverages gradient-based influence estimation.
+Given a small set of validation examples per task, LESS computes the influence of each training example by measuring the weighted cosine similarity between their LoRA gradients across multiple warmup checkpoints.
+It then aggregates these influence scores by averaging over validation examples within each task, followed by taking the maximum across tasks to obtain a scalar utility score per training example.
+Training examples are selected greedily based on these scores.
+In our experiments, we use the same base model for both selection and training, and follow the original LESS setup: 5% warmup training for 4 epochs and a gradient projection dimension of 8192.
+C.3 NV-Embed
+Embedding-based data selection as detailed in (Ivison et al., 2025) is a supervised data selection method that ranks training examples by computing cosine similarity between their embeddings and those of validation examples.
+Unlike model-aware methods like LESS, embedding-based data selection is model-agnostic: it relies on fixed, pretrained embedding models (in our case, we used NV-embed-v2, the state-of-the-art embedding model) rather than the target model.
+Instead of aggregating similarity scores into a single utility value per training example, embedding-based data selection uses a round-robin strategy that iteratively selects the highest-scoring example for each validation instance, ensuring diverse coverage across tasks.
+We follow the original setup from Ivison et al.
+(2025) in our experiments.
+Appendix D Further Training Details
 We train our models on a 4-GPU Nvidia-GH200 node, with batch size 256 and micro batch size 2.
-Appendix D Further Ablations on UltraInteract.
+Appendix E Further Ablations on UltraInteract.
+See Tables 11 and 4.
 Data Full UI Closest-1 Random-1
@@
 Table 12: The detailed comparison across benchmarks for self-distillation discussed in Section 5.5
-Appendix E Additional Related Works On Model Dependent Data Selection Approaches
-E.1 0. Notations
+Appendix F Additional Related Works On Model Dependent Data Selection Approaches
+Model-dependent data selection methods leverage internal signals from a target model—such as gradients, embeddings, or log-probabilities—to identify training examples that are most useful for fine-tuning.
+These approaches have led to strong empirical results across various settings.
+However, many of them involve substantial computational costs, such as repeated gradient computations or auxiliary model training, which can limit their scalability.
+We discuss these approaches and the costs they incur in this section.
+F.1 Notations
   1. 1.
@@
 Since the two parts above constitute the majority of computation for almost all the data selection approaches, we omit the other cost and only focus on these two.
-1. TLDR: The Final Table
+F.2 TLDR: The Final Table
 For GRAPE, we assume that in the training dataset $D$, various responses to the same instruction are already available, thus no additional cost is incurred in the Response Collection step of GRAPE.
@@
 Additional Training Per-Sample Feature Computation
-GRAPE (ours) 00 $NF_{\theta}$
+GRAPE (ours) 0 $NF_{\theta}$
 Gradient-based influence (LESS) $C(\theta_{\text{lora}},D_{\text{warmup}},T)$ $3T\cdot NF_{\theta}$
-In-run gradient-based influence $C(\theta,D,1)$ 00
+In-run gradient-based influence $C(\theta,D,1)$ 0
 Gradient matching $C(\theta_{\text{lora}},D_{\text{warmup}},T)$ $3T\cdot NF_{\theta}$
 Gradient norm $m\cdot C(\theta,D,1)$ $3m\cdot NF_{\theta}$
-Embedding-based 00 $NF_{\theta}$
-Simple uncertainty indicators 00 $NF_{\theta}$
+Embedding-based 0 $NF_{\theta}$
+Simple uncertainty indicators 0 $NF_{\theta}$
 Perplexity $C(\theta_{\text{ref}},D_{\text{ref}},1)$ $NF_{\theta_{\text{ref}}}$
@@
 Table 13: Computational cost comparison of data selection methods.
-2. Gradient-based Methods
+F.3 Gradient-based Methods
 Gradients have long been an important source of information for training data selection, as they directly affect the whole optimization process of language models.
@@
     Gradient norm
-2.1 Gradient-based Influence
+F.3.1 Gradient-based Influence
 Gradient-based influence computes the pairwise influence scores between each pair of training and validation examples.
 Training data with the highest influence are selected, as training on them leads to the theoretically largest decrease in model loss on validation data.
-LESS (Xia et al., 2024) formulates the pairwise influence scores as the cosine similarity between the gradients of training and validation data, and computes these gradient features using the following two steps:
+LESS Xia et al.
+(2024) formulates the pairwise influence scores as the cosine similarity between the gradients of training and validation data, and computes these gradient features using the following two steps:
   1. 1.
@@
   • Per-sample gradient: 0.
-2.2 Gradient Matching
+F.3.2 Gradient Matching
 Gradient matching also requires per-sample gradients, but utilizes their information in a different way.
 It performs clustering based on these gradient features to group similar data, and then applies an iterative greedy selection algorithm.
-In order to scale to LLM-level gradient computation and clustering, TAGCOS (TAGCOS) completely follows the warmup training and gradient computation pipeline of LESS (LESS).
-As the computational bottleneck here is still the gradient computation instead of clustering or iterative selection, (TAGCOS) also shares the same computational cost as (Xia et al., 2024):
+In order to scale to LLM-level gradient computation and clustering, TAGCOS Zhang et al.
+(2024b) completely follows the warmup training and gradient computation pipeline of LESS Xia et al.
+(2024).
+As the computational bottleneck here is still the gradient computation instead of clustering or iterative selection, Zhang et al.
+(2024b) also shares the same computational cost as Xia et al.
+(2024):
   • Additional training: $C(\theta_{\text{lora}},D_{\text{warmup}},T)$.
   • Per-sample gradient for each checkpoint: $NT\cdot 3F_{\theta}=3T\cdot NF_{\theta}$.
-2.3 Gradient Norm
+F.3.3 Gradient Norm
 The $L_{2}$-norms of gradient vectors can also serve as effective indicators for data selection.
-(Paul et al., 2023) proposes GraNd, which obtains a utility score for each training point based on its gradient norm early in the training.
+Paul et al.
+(2023) proposes GraNd, which obtains a utility score for each training point based on its gradient norm early in the training.
 More specifically, it starts from $m$ different model weight initializations, trains each model on the whole dataset to obtain per-sample gradient norms, and finally averages the $m$ gradient norms for each training point to obtain the final GraNd score.
@@
   • Per-sample gradient for each weight initialization: $Nm\cdot 3F_{\theta}=3m\cdot NF_{\theta}$.
-3. Embedding-based Methods
+F.4 Embedding-based Methods
 Embedding-based methods project the whole training set into an embedding space to quantify the information of each data point and their interactions.
 For model-based embedding-based selection methods, the embeddings are usually computed by the final model $\theta$ to align with its preference.
-Under a supervised data selection setup where validation data representing target task distributions are available, Representation-based Data Selection (RDS; (Rubin et al., 2022; Hanawa et al., 2021)) computes the embedding similarity between training and validation data, and selects training points that are most similar to the target distribution in the embedding space.
-For an unsupervised setup where only the embeddings of training data are accessible, geometry-based coreset sampling methods are widely used (Qin et al., 2024).
+Under a supervised data selection setup where validation data representing target task distributions are available, Representation-based Data Selection (RDS; Rubin et al.
+(2022); Hanawa et al.
+(2021)) computes the embedding similarity between training and validation data, and selects training points that are most similar to the target distribution in the embedding space.
+For an unsupervised setup where only the embeddings of training data are accessible, geometry-based coreset sampling methods are widely used Qin et al.
+(2024).
 Grounded on the intuition that close samples in the embedding space often share similar properties, a diverse subset can be obtained by controlling the minimum distance between any two selected data points.
-Among them, using K-center greedy sampling to select embedding-based facility locations has been proven especially effective for instruction fine-tuning of LLMs (Bhatt et al., 2024a).
+Among them, using K-center greedy sampling to select embedding-based facility locations has been proven especially effective for instruction fine-tuning of LLMs Bhatt et al.
+(2024).
 These embedding-based approaches share similar computational costs: they do not need any additional model training and can directly extract useful per-sample embeddings using the last-layer hidden states of the pretrained final model $\theta$.
@@
   • Per-sample embedding computation: $NF_{\theta}$.
-4. LogProb-based Methods
+F.5 LogProb-based Methods
 LogProb-based methods also directly utilize the target LLM to evaluate the utility of each training data point.
 4.1 Simple Uncertainty-based Indicators
-Some simple model-based indicators inspired by the notion of uncertainty have been shown effective for a long time and recently extended to data selection for LLM instruction tuning (Marion et al., 2023b; Bhatt et al., 2024b).
-(Bhatt et al., 2024b) demonstrates the effectiveness of various indicators including mean entropy, least confidence, mean margin, etc.
+Some simple model-based indicators inspired by the notion of uncertainty have been shown effective for a long time and recently extended to data selection for LLM instruction tuning (Marion et al., 2023b; Bhatt et al., 2024).
+Bhatt et al.
+(2024) demonstrates the effectiveness of various indicators including mean entropy, least confidence, mean margin, etc.
 These simple indicators do not require additional training and can also be directly obtained with the pretrained final model $\theta$.
@@
   • Per-sample per-token logits computation: $NF_{\theta}$.
-4.2 Perplexity (PPL)
+F.5.1 Perplexity (PPL)
 PPL is also a long-standing data selector and has been shown effective for LLM-scale data selection.
 Typically, a split of the training dataset, $D_{\text{ref}}$, is needed to train $\theta_{\text{ref}}$, a reference model that will be used to compute PPL for the whole training set.
-A common approach is to use the final model $\theta$ as the reference model $\theta_{\text{ref}}$ to ensure the alignment in PPL patterns (Marion et al., 2023a), but prior work (Ankner et al., 2024) also shows that a reference model much smaller than the final model can also be an effective PPL-based data selector.
+A common approach is to use the final model $\theta$ as the reference model $\theta_{\text{ref}}$ to ensure the alignment in PPL patterns Marion et al.
+(2023a), but prior work Ankner et al.
+(2024) also shows that a reference model much smaller than the final model can also be an effective PPL-based data selector.
 The computational cost for PPL-based selection is shown below:
@@
   • Per-sample PPL computation: $NF_{\theta_{\text{ref}}}$.
-4.3 Learnability
+F.5.2 Learnability
 In addition, learnability (Mindermann et al., 2022; Zhou et al., 2024a, b) is a more effective metric than pure uncertainty or PPL, as it excludes uncertain but unlearnable points (e.g., noisy or less task-relevant) by considering the decrease in per-sample loss before and after the model is fully trained.
@@
   • Per-sample learnability computation: $2\cdot NF_{\theta}$.
-4.4 Loss Trajectory
+F.5.3 Loss Trajectory
 Moreover, logprob-based methods can also obtain finer-grained information from the training dynamics of LLMs.
-S2L (Yang et al., 2024b) obtains a feature vector for each training point by collecting their training loss trajectories over $T$-epoch training on a small reference model $\theta_{\text{ref}}$, and then applies K-means clustering to equally sample data points from each trajectory cluster.
+S2L Yang et al.
+(2024b) obtains a feature vector for each training point by collecting their training loss trajectories over $T$-epoch training on a small reference model $\theta_{\text{ref}}$, and then applies K-means clustering to equally sample data points from each trajectory cluster.
 Prior work shows its superiority over other logprob-based indicators, but it also comes with significant computational cost:
   • Additional training: $C(\theta_{\text{ref}},D,T)$.
-Here the choice of $\theta_{\text{ref}}$ is especially important, as prior work (Xia et al., 2023) shows that reference models that come from the same model family as the final model tend to have similar loss trajectories of training data, so they can preserve more fidelity in their loss trajectory patterns.
+Here the choice of $\theta_{\text{ref}}$ is especially important, as prior work Xia et al.
+(2023) shows that reference models that come from the same model family as the final model tend to have similar loss trajectories of training data, so they can preserve more fidelity in their loss trajectory patterns.
   • Per-sample loss trajectory computation: $T\cdot NF_{\theta_{\text{ref}}}$.
 Note that $T$ here is typically much larger than that in gradient-based influence computation, so the computational cost of this gradient-free approach can be even higher than gradient-based methods.
+Appendix G Multi-Round GRAPE
+To evaluate whether GRAPE continues to provide benefits after an initial round of fine-tuning, we conducted a second-round experiment using LLaMA3.1-8B.
+In the first round, we selected the best responses per instruction using GRAPE and fine-tuned the model accordingly.
+Then, we re-applied GRAPE on the newly fine-tuned model to select a fresh set of responses and conducted another round of fine-tuning.
+This second iteration led to further performance improvements across multiple benchmarks, including AlpacaEval, WizardEval (both original and LeetCode), and real-world tasks such as MATH, MMLU, and BBH.
+Notably, GRAPE Round 2 improved average benchmark scores from 35.9% to 37.3%, demonstrating that the method remains effective and even compounding when iteratively applied.
+Method AE WR WR (LC) MATH MMLU BBH
+GRAPE Round 1 15.2 14.8 19.4 32.1 64.5 69.6
+GRAPE Round 2 17.6 19.6 18.9 33.2 64.5 70.0
+Table 14: Performance of LLaMA3.1-8B after two rounds of GRAPE fine-tuning.
+Appendix H Details Of Correlation Analysis
+To analyze how well different models align with the training distribution, we conducted a perplexity-based study on the Tulu-v3 training set.
+Specifically, we randomly sampled 1,000 instances from the training data and used a wide array of generator models to produce responses for each instruction.
+For each response, we computed its perplexity under each base model and ranked the responses by perplexity, with lower perplexity indicating better matching of distribution.
+We then identified the top-1 response per instance for each base model and visualized the overall rankings using a heatmap (Figure 4).